Notes on tameness Brian H. Bowditch Mathematics Institute, University of Warwick, Coventry, CV4 7AL, Great Britain. [First draft: June 2006. Revised: November 2009]

0. Introduction.

The aim of this paper is to give an account of the Tameness Theorem for complete hyperbolic 3-manifolds. Our presentation makes use of ideas from a number of different sources. We aim to give an account that is readily accessible from fairly standard geometrical and topological arguments, and adaptable to variable curvature.

Let M be a complete hyperbolic 3-manifold (without boundary) with π_{1}(M) finitely generated. Marden [Ma] asked if this implies that M is topologically finite, i.e. homeomorphic (or equivalently diffeomorphic) to the interior a compact manifold with boundary. This was proven in the “indecomposable” case by Bonahon [Bon2], and in general independently by Agol [Ag] and Calegari and Gabai [CalG]. This is generally referred to as the “Tameness Theorem”:

Theorem 0.1 : (Bonahon, Agol, Calegari, Gabai) Let M be a complete hyperbolic 3-manifold with π_{1}(M) finitely generated. Then M homeomorphic to the interior of a compact manifold with boundary.

(It is well known that such a compactification is unique up to homeomorphism.) In view of [Tu] it is easily seen that a 3-manifold is topologically finite if and only if

its orientable cover is. For this reason we will assume henceforth that all our 3-manifolds are orientable. However, all the theorems stated in this section remain valid in the nonorientable case. We also note that, using [MeS], Theorem 0.1 extends to orbifolds, as we explain in Section 6.6.

To set Theorem 0.1 in context, we note that (in contrast to the analogous statement for 2-manifolds) one certainly needs the geometrical hypothesis in Theorem 0.1. For example Whitehead [Wh] described a simply connected 3-manifold which is not topologically finite in the above sense (see also [H]). We do however have the following positive topological result due to Scott [Sc1,Sc2]:

Theorem 0.2 : (Scott) Let M be a (topological) 3-manifold with π_{1}(M) finitely generated. Then there is a compact submanifold of M whose inclusion into M induces an isomorphism of fundamental groups.

Such a compact submanifold is referred to as a Scott core.

This is a key ingredient in the proof of the Tameness Theorem. It effectively reduces it to asserting that the ends of M are just topological products.

Notes on tameness

In fact, we are only interested here in irreducible 3-manifolds (that is, where every embedded 2-sphere bounds a ball). This always holds for a hyperbolic 3-manifold. Such a manifold is aspherical and it follows that we can obtain a homotopy equivalence in Scott’s theorem (after capping off any 2-spheres in the boundary by 3-balls). Moreover, irreducibility allows us to bypass discussion of the Poincaré conjecture, as proven by Perelman, since no such manifold can contain a fake 3-ball.

The Tameness Theorem is related to another well known result, namely Ahlfors’s Finiteness Theorem [Ah]. This is usually phrased analytcally in terms of Riemann surfaces. However (in the torsion-free case) it is easily seen to be equivalent to the following geometrical statement. Let M be a hyperbolic 3-manifold, which we assume non-elementary. (This rules out certain trivial cases where π_{1}(M) is trivial, Z or Z⊕Z.) The “convex core” of M is essentially the unique smallest submanifold which carries the whole of π_{1}(M) and has convex boundary. (It follows that its inclusion into M in a homotopy equivalence.) We say “essentially” since it might degenerate to a totally geodesic surface in M. In any case, for any t > 0, the metric t-neighbourhood of the convex core is a C^{1}-submanifold. (Unlike the Scott core, the convex core need not be compact.) The boundary, F_{t}, of this neighbourhood is an embedded C^{1}-surface (possibly empty or disconnected). As we note in Section 8, Ahlfors’s Finiteness Theorem (in the torsion-free case) is equivalent to:

Theorem 0.3 : Let M be a non-elementary complete hyperbolic 3-manifold with π_{1}(M) finitely generated. Let t > 0, and let F_{t} be the boundary of the t-neighbourhood of the convex core of M. Then F_{t} has finite area.

In fact, we can make similar definitions when M is complete riemannian 3-manifold of pinched negative curvature. We have versions of both tameness and the Ahlfors Finiteness Theorem respectively generalising Theorems 0.1 and 0.3:

Theorem 0.4 : Let M be a complete riemannian 3-manifold of pinched negative curvature with π_{1}(M) finitely generated. Then M is homeomorphic to the interior of a compact manifold with boundary.

Theorem 0.5 : Let M be a complete riemannian 3-manifold of pinched negative curvature with π_{1}(M) finitely generated. Let t > 0, and let F_{t} be the boundary of the t-neighbourhood of the convex core of M. Then F_{t} has finite area.

These are proven in Sections 7 and 8 respectively. As far as I know, these have not been made explicit before, though it seems to have been folklore that tameness techniques could be applied to give results of this type in variable curvature.

Since the work of Agol and Calegari and Gabai, other accounts of tameness have been given by Choi [Ch] and Soma [Som] (in the case without cusps). Our exposition is inspired by that of Soma, which gives another perspective on various constructions of [CalG]. The strategy we follow is broadly similar. However we phrase things differently, so as to remove the dependence on the end reduction theory [BrinT,My] and we include an independent argument for the relevant part of Souto’s result [Sou]. In this way, almost all our arguments, are built from first principles from results essentially known prior to Thurston’s work [Th1]. Apart, that is, from one critical appeal to the hyperbolisation theorem for atoroidal Haken manifolds [Th2,O,Ka], which is here used to deduce a purely topological statement (Theorem 2.8.1). There are a couple of places where the argument could be shortened a little by appeal to other more sophisticated results or machinery, as we point out in the relevant places.

Originally, the term “tameness” was used by Thurston [Th1] to refer to a certain geometric property of M. He showed that this implies topological finiteness. Subsequently Canary [Can1] showed that tameness was, in fact, equivelant to topological finiteness. Since then, the two terms have commonly been regarded as synonymous, though many of the consequences of tameness follow via Thurston’s geometrical interpretation.

One consequence of tameness is the Ahlfors measure conjecture, which states that the limit set of a finitely generated kleinian group has either zero or full Lebesgue measure (see [Can1]). Tameness is also a critical step in the classification of finitely generated kleinian groups. Other key ingredients in this classification are Thurston’s Ending Lamination Conjecture, [Mi,BrocCM] (see also [Bow2,Bow3]), as well as a description of those end invariants which are realisable (see for example [KlS,NS]). Furthermore, in combination with various other results, it leads to a proof of the density conjecture of Bers, Sullivan and Thurston, namely that geometrically finite kleinian groups are dense among finitely generated ones. Arguments to deal with many (but not all) cases have been given by Bromberg and Brock [Brom,BrocB].

A key notion we use is that of a “polyhedral” surface, as in [Som]. The basic idea is that, appropriately interpreted, many of the essential constructions of “pleated surfaces” in hyperbolic 3-manifolds adapt to a more general context. Such surfaces have also been used, for example, in [Bon2] (see Lemme 1.7) and [Can2] (Section 3 thereof), where they are termed “simplicially hyperbolic”. They also feature in [Sou].

As seems to be the tradition, we simplify the exposition by dealing first with the case where there are no parabolics (Section 5), and later discuss how this generalises when there are parabolics (Section 6). We go on to discuss variable curvature (Section 7) and Ahlfors’s Finiteness Theorem (Section 8).

Some of the material for this paper was worked out while visiting the Université Paul Sabatier in Toulouse. I am grateful to that institution for its support. It was mostly written at the University of Southampton, and revised at the University of Warwick. I thank Dick Canary for his comments on an earlier draft of this paper.

1. Outline of the proof.

In this section, we give a brief informal outline of the proof of tameness. We restrict the discussion here to the case without cusps. The adaptations necessary when there are cusps will be discussed in Section 6.

Let M be a complete hyperbolic 3-manifold with π_{1}(M) finitely generated. Using Scott’s theorem, it is enough to show that each end, e, of M is topologically finite; in other words, it has a neighbourhood that is homeomorphic to a closed surface times [0, ∞). As discussed in Section 2.4, e has a certain “genus”, g = genus(e), associated to it, which is determined a-priori. Using Waldhausen’s cobordism theorem, we can show that e is topologically finite if we can find a sequence of embedded surfaces of genus g going out the end, which separate the end from a fixed compact core the manifold and which are all homotopic to each other in the complement of the core — see Corollary 2.5.2. In fact we can make do with a sequence of maps of a fixed closed surface of genus g into M which homologically separate the end e from a core (since such maps can be replaced by embeddings using [FrHS] as described in Section 2.6). As discussed in Section 2.4, after passing to a cover, we can assume that all of π_{1}(M) is carried by (any neighbourhood of) the end, e.

So far, the discussion has been topological. We now assume that M is hyperbolic without cusps. One can split into two cases. If e is “geometrically finite”, then topological finiteness follows easily. We therefore assume that e is “degenerate”, i.e. not geometrically finite. We construct a compact polyhedral subset Φ = ϕ(K), of M by taking a suitable finite simplicial 2-complex, K, and mapping it into M by a piecewise geodesic map ϕ : K ⟶ M, where ϕ is homotopy equivalence. In fact, we can extend this to a larger complex, ϕ : K_{α} ⟶ M, to give a larger polyhedral subset Φ_{α} ⊆ M. Since e is assumed degenerate, we can arrange that Φ_{α} enters arbitrarily far into e (see Lemma 5.2). Its purpose will be to act as a barrier to the homotopies we want to perform. We can arrange that the completion of the complements of Φ and Φ_{α} are locally CAT(−1) spaces.

Now let F be a closed embedded surface in M \ Φ_{α} separating e from Φ_{α}. If it happens that its genus equals g = genus(e) then its inclusion in M \Φ_{α} will be π_{1}-injective (Corollary 2.4.5), and we can homotope it to a “balanced” polyhedral map in (the metric completion of) M \ Φ_{α}. Via a Gauss-Bonnet argument, we get geometric control on the diameter of the image of this map. We end up with a map of Σ into M which homologically separates e from Φ and (since Φ_{α} enters deeply into e) stays a long way from Φ. If we could find such an F for every such Φ_{α}, then we get a sequence of surfaces, which we can replace by embeddings using [FrHS], as described in Section 2.6. One can then give an argument (cf. [Sou]) to show that these lie in finitely many homotopy classes in M \ Φ (see Lemma 5.3). We can therefore assume they all lie in the same homotopy class and we are done by Waldhausen as described in the previous paragraph.

The problem is that, a-priori, there is no reason to suppose that we can choose F with the right genus. In general its genus will be greater than g (Lemma 2.4.4). Using Dehn’s Lemma, we can certainly assume that its inclusion into M \Φ_{α} is π_{1}-injective. The trick now (at least in the compressible case) is to consider the component, W, of M \ F, containing Φ_{α}. Let X be the cover of W corresponding to the image of π_{1}(K_{α}) ≅ π_{1}(M) in W. We lift ϕ to a map ϕ̃: K_{α} ⟶ M and set Φ̃_{α} = ϕ̃(K_{α}) ⊆ W, and Φ̃ = ϕ̃(K) ⊆ Φ̃_{α}. Using an inductive hypothesis on genus, and Theorem 2.8.1, one can show that X is topologically finite. It has an “outer” end which corresponds to e. Removing a collar neighbourhood of this outer end, we obtain a manifold P ⊆ X, with boundary ∂P, so that X \ P is just a product ∂P × [0, ∞). We can assume Φ̃_{α} ⊆ P. Moreover, it turns out that ∂P has the correct genus g. This surface, ∂P, can now play a similar role to F as in the previous paragraph.

It would be nice to homotope ∂P in P \ Φ̃_{α} to a balanced polyhedral map there, and then map down to M to give us a surface homologically separating e from Φ, and with controlled diameter. There are a couple of complications however.

First, the homotopy might get snagged on ∂P, over which we have no geometric control. To get around this, we first embed P via a homotopy equivalence larger manifold, Z, with polyhedral boundary. This space Z will be locally CAT(−1), and the natural map of P to M extends to Z. We can then carry out our straightening homotopy in Z instead. To constrain the amount this homotopy moves the surface, we consider its homological intersection with a certain ray, τ, going out the end.

Another complication is that, although our homotopy is disjoint from Φ̃_{α} in Z, its projection to M may sweep through Φ_{α}, or indeed through Φ. We therefore need some other argument to show that the polyhedral surface we end up with does indeed separate e from Φ. The intuitive reason for this is that the homotopy does not meet the preimage, Ψ ∩ int Z, of Φ in int Z. In fact, Ψ ∩ int Z has only one compact component, and this maps homeomorphically to Φ (cf. Lemma 4.6). The homotopy cannot sweep through any non-compact component. We shall approach this more formally, in the form of Lemmas 4.7 and 4.8.

We now choose larger and larger sets Φ_{α}, to give us our sequence of surfaces, and we are done (in the compressible case without cusps) using [FrHS], Lemma 5.3 and Waldhausen as described above.

A slight variation on the above is called for in the case where the end e is incompressible. Most of the essential core of the above argument is to be found in Sections 4 and 5.

2. Topological constructions.

In this section, we review or modify various standard 3-manifold constructions. The arguments are purely topological, apart from one appeal to Thurston’s hyperbolisation theorem in Section 2.8, albeit to prove a purely topological statement. We divide the discussion into a number of largely independent topics. We confine our discussion to the “non-relative” case — sufficient to deal with the case without cusps. Some generalisations to the relative case will be discussed in Section 6.

Convention : We assume all 3-manifolds to be aspherical (every 2-sphere bounds a ball) and orientable.

The former assumption holds in any hyperbolic 3-manifold. The latter will simplify the exposition, and as observed in the introduction, is justified by [Tu].

We will reserve the letter M to refer to 3-manifolds with empty boundary. A “complete hyperbolic 3-manifold” is assumed to have empty boundary. In general, topological 3manifolds allowed to have non-empty boundary.

2.1. Basic notions.

We refer to [H] for the general theory. In particular, we note the following.

Let P be a 3-manifold with (possibly empty) boundary ∂P.

Definition : A compressing disc is a properly embedded disc, D ⊆ P, with ∂D = D ∩ ∂P an essential closed curve in ∂P.

We have “Dehn’s Lemma” due to Papakyriakopoulos:

Lemma 2.1.1 : [Papakyriakopoulos] If F ⊆ ∂P is any subsurface of ∂P with π_{1}(F) ⟶ π_{1}(P) not injective, then there is a compressing disc, D ⊆ P with ∂D ⊆ F.

We also have the following special case of work of Stallings. It can be proven directly by the argument given for example in “Stage 3” of [Sc2]:

Lemma 2.1.2 : If P is compact and π_{1}(P) is a non-trivial free product, then P has a compressing disc.

We also note from Papakyriakopoulos’s Sphere Theorem that (since P is assumed irreducible) π_{2}(P) = 0.

We can also define compressing discs for closed surfaces in P:

Definition : If S ⊆ P is an embedded 2-sided closed surface, a compressing disc for S is an embedded disc, D ⊆ P, with ∂D = D ∩ S an essential closed curve in S. We say that S is incompressible if it admits no compressing disc.

Using Dehn’s lemma (applied to P cut along S) one can show that this is equivalent to saying that π_{1}(S) ⟶ π_{1}(P) is injective.

Definition : We say that P is atoroidal if every incompressible torus in P is boundary parallel (i.e. can be homotoped into a boundary component).

2.2. Compression bodies.

Let H be a handlebody and let Σ = ∂H. If we imagine H as embedded unknotted in R^{3}, then a compression body, P, is obtained by removing the interior of a (possibly empty) set of unknotted positive genus handlebodies H_{1}, H_{2}, …, H_{n} from the interior of H, such that the H_{i} are separated by a collection of disjoint compressing discs in H. (Here “positive genus” means not a 3-ball.) Write Σ_{i} = ∂H_{i}. Thus, ∂P = Σ ∪ Σ_{1} ∪ ⋯ Σ_{n}. We refer to ∂_{0}P = Σ as the outer boundary, and to the Σ_{i} as the inner boundary components. Each inner boundary component is incompressible.

(One can equivalently define compression bodies inductively. Start with a disjoint collection of handlebodies and products of surfaces with the interval, and then glue them together along discs in their respective outer boundaries so as to form a connected manifold.)

Let g = genus(Σ), g_{i} = genus(Σ_{i}) and g_{0} = g−∑_{i=1}^{n} g_{i} ≥ 0. We refer to (g_{0}, g_{1}, …, g_{n}) as the type of P. It determines P up to homeomorphism. Note that special cases are those of a handlebody (type (g)) or a product Σ × [0, 1] (type (0, g)). In the latter case, there is no natural distinction between the inner and outer boundaries. In all other cases, the outer boundary is the unique maximal genus boundary component.

Note that

π_{1}(P) = F_{g_{0}} ∗ π_{1}(Σ_{1}) ∗ ⋯ ∗ π_{1}(Σ_{n})

where F_{p} denotes the free group of rank p. Recall that the non-cyclic factors in a maximal free product decomposition are unique up to conjugation. We refer to a group of this sort as a compression group of type (g_{0}, g_{1}, …, g_{n}). (This is termed a “free/surface group” in [CalG].) We refer to g = g_{0} + ⋯ + g_{n} as its genus. One can show (cf. [CalG] Lemma 5.6):

Lemma 2.2.1 : If P is a compact (orientable and irreducible) 3-manifold with ∂P ≠ ∅, with π_{1}(P) a compression group, and with each surface factor is conjugate to the fundamental group of boundary component, then P is a compression body.

Note that the genus of the group is the genus of the outer boundary component. The clause about surface groups factors can be omitted if π_{1}(P) is itself a surface group. (The only purpose of the hypothesis ∂P ≠ ∅ is to rule out a simply connected closed 3-manifold.)

Proof : (Sketch) We refer to the boundary components homotopic to the surface factors as “inner” boundary components. From the hypotheses these are all π_{1}-injective. (A priori, there might be any finite number of non-inner boundary components.) If π_{1}(P) is trivial then each component of ∂P is a 2-sphere and so P is a 3-ball. If π_{1}(P) is a surface group, then P is a closed surface times an interval [H]. Otherwise P splits as a non-trivial free product, and so Lemma 2.1.2 gives us a compressing disc D ⊆ P. Thus ∂D lies in a non-inner boundary component of ∂P. We cut P along D, and verify that the hypotheses also hold for the resulting manifold, or pair of manifolds. By induction on the genus of the compression group, we can suppose these pieces are compression bodies. The original manifold P therefore consists either of a compression body with a handle attached to the outer boundary, or two compression bodies connected along a disc in their respective outer boundaries. Thus, P is itself a compression body. ♢

We also have the following (a stronger version of which can be found in [BrinJS]).

Lemma 2.2.2 : If P is a compact 3-manifold and Σ ⊆ ∂P is a boundary component such that the induced map π_{1}(Σ) ⟶ π_{1}(P) surjects, then P is a compression body with outer boundary ∂_{0}P = Σ.

Proof : (Sketch) The argument is similar to that of Lemma 2.2.1, This time, the outer boundary is determined a-priori by the hypotheses, and we use Dehn’s Lemma (2.1.1) in place of Lemma 2.1.2 to find a compressing disc. ♢

Let P be a compression body of type (g_{0}, g_{1}, …, g_{n}). Abelianising, we see that H_{1}(P) (with Z_{2} coefficients) decomposes as H_{1}(P) ≡ Z_{2}^{g_{0}} ⊕ ⨁_{i} H_{1}(Σ_{i}), where H_{1}(Σ_{i}) = Z_{2}^{2g_{i}}.

We also note that H_{2}(P) ≡ ⨁_{i} H_{2}(Σ_{i}). In particular, H_{2}(∂P) ⟶ H_{2}(P) surjects, and so H_{2}(P, ∂P) = 0. Thus any closed surface in P separates.

2.3. Cores.

Here we shall take homology with Z_{2} coefficients, though the discussion applies more generally.

Let M be a non-compact 3-manifold (without boundary). An end of M is isolated if it is seperated from all other ends by a compact subset of M. (If M has only finitely many ends then all of them are isolated.) Any isolated end, e, of M has an associated (possibly trivial) second homology class, h(e) ∈ H_{2}(M) — induced by any surface that separates the end.

If P ⊆ M is a compact connected submanifold with H_{2}(P) ⟶ H_{2}(M) surjective, then P separates the ends of M. In particular, M has finitely many ends (all isolated). If H_{2}(P) ⟶ H_{2}(M) is also injective, then each component of M \ int P is non-compact. We see that there is a bijection between the ends of M and the components of M \ int P. If H_{1}(P) → H_{1}(M) is surjective, then each such component has connected boundary, and so there is a bijection between the ends of M and the boundary components of P. In particular, this applies if P ↪ M is a homotopy equivalence, or more generally, if P carries all of H_{1} and H_{2} and no component of M \ int P is compact.

We can identify the image of H_{2}(∂P) in H_{2}(P) ≡ H_{2}(M) with the span of h(e) as e varies over the set of ends. If H_{2}(P) = 0, there is precisely one relation arising from the fact that H_{3}(P, ∂P) = Z_{2} is the kernel of H_{2}(∂P) ⟶ H_{2}(P).

We recall again, the key result (Theorem 0.2):

Theorem 2.3.1 : (Scott [Sc1,Sc2]) Suppose M is a 3-manifold with π_{1}(M) finitely generated. Then there is a compact submanifold P whose inclusion into M is a homotopy equivalence. ♢

We refer to P as a Scott core of M.

2.4. Ends.

Let M be a 3-manifold with π_{1}(M) finitely generated, and let e be an end of M. Let E ⊆ M be a submanifold with ∂E compact, containing the end e. This will be a “neighbourhood” of the end e. Given a subset, Q ⊆ M, we write G(Q) ≤ π_{1}(M) for the image of π_{1}(Q) in π_{1}(M).

Definition : We say that E is full if ∂E is connected and we can homotope M into M \E.

(For example, this holds if M\E contains a Scott core.) In particular, π_{1}(M\E) ⟶ π_{1}(M) is surjective (i.e. G(M \ E) = π_{1}(M)). It also follows that E contains no other end of M. We have maps π_{1}(∂E) ⟶ π_{1}(E) ⟶ π_{1}(M). Since π_{1}(M) = π_{1}(M \ E) ∗_{G(∂E)} G(E), this splitting must be trivial, and so G(∂E) = G(E) (with respect to a basepoint in ∂E). In particular, G(E) is finitely generated.

Let N = N(E) be the cover of M corresponding to G(E). Now E lifts to a one-ended submanifold, Ê ⊆ N which carries all of π_{1}(N). This is a neighbourhood of an end, ê, of N. In fact, the homotopy of M into M \ E lifts to a homotopy of N into N \ E, and so Ê is a full neighbourhood of ê in N.

We claim that N(e) is independent of the choice of full neighbourhood of e:

Lemma 2.4.1 : Suppose E and E′ are full neighbourhoods of e in M. Taking any basepoint in the unbounded component of E ∩ E′, the subgroups G(E) and G(E′) of π_{1}(M) are equal.

Proof : Since there is a base of full neighbourhoods of e, we can assume that E ⊆ E′. Clearly G(E) ⊆ G(E′). Let λ : N ⟶ M be the cover of M corresponding to G(E). As observed above, E lifts to a (full) neighbourhood, Ê, of an end ê of N. Since E′ is full in M, we can homotope ∂E into M \ E′ in M. We can lift this to a homotopy in N, starting at ∂Ê and finishing with a map into N, with image Q say, disjoint from Ê. Let R ⊆ N be the set of point to which this homotopy maps with degree 1. Thus, R is compact, and ∂Ê ⊆ ∂R ⊆ Q ∪ ∂Ê. Moreover, R lies on the opposite side of ∂Ê to ê. Now λ(Q) ⊆ M \ Ê, and so ∂E′ ⊆ λ(R). Since λ is covering map, it follows that λ|(R ∩ λ^{−1}(∂E′)) is a homeomorphism. In other words we can lift ∂E′ to N. It follows that G(E′) = G(∂E′) ⊆ G(E), so G(E′) = G(E) as claimed. ♢

In particular, we see that the isomorphism type of G(E) depends only on e, and we denote it by G(e) — the end group. The cover N(e) = N(E) only depends on E. We will denote the corresponding end of N(e) also by e.

Let N = N(e). Let P be a Scott core in N, and let E be the component of N \ int P containing e. Thus, E is a full neighbourhood, and so G(E) = G(e) = π_{1}(N). Thus π_{1}(∂E) ⟶ π_{1}(N) is surjective. But this factors through π_{1}(P) ⟶ π_{1}(N) which is an isomorphism, and so π_{1}(∂E) ⟶ π_{1}(P) is also surjective. By Lemma 2.2.2, P is a compression body with outer boundary ∂_{0}P = ∂E. Thus, G(e) ≅ π_{1}(P) is a compression group. We have shown:

Lemma 2.4.2 : The end group, G(e), of any end e of M is a compression group. ♢

Its type, (g_{0}, g_{1}, …, g_{n}) is well defined, and we refer to g = g_{0} + ⋯ + g_{n} as the genus of the end e of M (or of N). We denote it by genus(e).

Remark : One can go on to show that genus(e) is equal to the genus of the corresponding boundary component of any Scott core of M, though we won’t be needing this fact here.

Definition : We say that an end, e, of M is incompressible if it has type (0, g) for some g ∈ N (the genus of e).

In other words, the cover N(e), has the homotopy type of a closed surface, Σ, of genus g. Suppose S ⊆ N(e) is any surface separating the ends of N(e). We get a degree-1 map of S to Σ. If genus(S) = g, then this is a homotopy equivalence (using the Hopf property of π_{1}(Σ)). Suppose E is the neighbourhood of e with ∂E = S. Then the inclusion of S = ∂E into E is also a homotopy equivalence (for example, via van-Kampen’s theorem and using asphericity). It follows that E deformation retracts onto ∂E, and that E is a full neigbourhood of e in N(e). Note that (from the earlier discussion) such a surface, S, always exists in N(e).

The following is now easily verified:

Lemma 2.4.3 : If e has a neighbourhood, E, in M with ∂E incompressible M, then e is incompressible, and E deformation retracts onto ∂E. ♢

Remark : In fact, such a neighbourhood will always exist in M if e is incompressible — take a complementary component of a Scott core — though we won’t need this fact here.

We need the following result about surfaces separating an end, e, of M.

Lemma 2.4.4 : Suppose that E is a full neighbourhood of the end e. Then genus(∂E) ≥ genus(e).

Proof : We can pass to the cover, N = N(e), of M corresponding to e. As observed earlier, E is also full in N. Let e_{1}, …, e_{n} be the set of other “inner” ends of N. (This set may be empty.) These ends are all incompressible in N. (Take a Scott core of N and apply Lemma 2.4.3.) We now take a maximal compression of ∂E in N \ int E, so as to give us a compression body Q ⊆ N \int E with outer boundary ∂_{0}Q = ∂E, and incompressible inner boundary components, S_{1}, ⋯, S_{n}. (Take a maximal collection, D_{1}, …, D_{k}, of disjoint compressing discs for ∂E in N \ int E, thicken up ∂E ∪ ⋃_{i} D_{i} in N \ E, and cap off with a 3-ball any 2-sphere boundary components that arise.)

Recall that π_{1}(N) is a compression group, F_{g_{0}} ∗ (∗_{i}π_{1}(Σ_{i})), where Σ_{i} is a surface of genus g_{i}, embedded in N, and separating the end e_{i}. Now π_{1}(S_{i}) does not split as a free product, and is thus conjugate in π_{1}(N) to one of the surface groups π_{1}(Σ_{j}). Thus, S_{i} is homotopic in N to Σ_{j} and so separates the end e_{j}. No two S_{i} can correspond to the same end e_{j}, since they would then bound a compact region. Moreover, we see that the S_{i} must account for all of the inner ends of N (since each such end is contained in some component of N \ int Q bounded by some S_{i}). We can thus assume that S_{i} separates the end e_{i}. Let E_{i} be the component of N \ int Q containing e_{i}, so that ∂E_{i} = S_{i}. Since e_{i} is incompressible, E_{i} deformation retracts onto ∂E_{i}. Thus π_{1}(Q) ⟶ π_{1}(N \ int E) is surjective. Since E is full, π_{1}(N \ E) ⟶ π_{1}(N) is surjective, and so π_{1}(Q) ⟶ π_{1}(N) is also surjective. Now π_{1}(Q) is a compression group of type (h, g_{1}, …, g_{n}) for some h ≥ 0, and we have a surjective map F_{h} ∗ (∗_{i}π_{1}(S_{i})) ⟶ F_{g_{0}} ∗ (∗_{i}π_{1}(Σ_{i})) which sends π_{1}(S_{i}) bijectively to π_{1}(Σ_{i}). Quotienting out by the normal closures of the surface factors, we get a surjective map between free groups, F_{h} ⟶ F_{g_{0}}. It follows that h ≥ g_{0}.

But genus(e) = g_{0}+g_{1}+⋯+g_{n} and genus(∂E) = h+g_{1}+⋯+g_{n}, and so genus(∂E) ≥ genus(e) as required. ♢

Corollary 2.4.5 : Suppose that P ⊆ M is a closed subset and that S ⊆ M \ P is a connected surface separating P from e. Suppose that the component of M \S containing e can be homotoped into P. Then genus(S) ≥ genus(e). Moreover, if genus(S) = genus(e), then S is π_{1}-injective in M \ P.

Proof : Note that S bounds a full neighbourhood, E, of e in M, and so the inequality follows immediately from Lemma 2.4.4. For the second statement, note that if the conclusion fails, then we could surger S along a compressing disc in M \ P to give surfaces of smaller genus, one of which must also separate e from P giving a contradiction. ♢

2.5. Waldhausen’s cobordism theorem.

Theorem 2.5.1 : (Waldhausen [Wa]) Suppose S_{1}, S_{2} ⊆ M are disjoint embedded π_{1}-injective surfaces of positive genus in a 3-manifold M, and suppose that S_{1} can be homotoped in M to S_{2}. Then there is an embedded submanifold R ⊆ M, homeomorphic to S_{1} × [0, 1], with ∂R = S_{1} ∪ S_{2}. ♢

Definition : We say that an end, e, of M is topologically finite if it has a neighbourhood homeomorphic to Σ × [0, ∞) for some surface Σ.

Corollary 2.5.2 : Suppose that P ⊆ M is closed, and (S_{i})_{i} is a sequence of surfaces in M \ P separating P from an end e of M, and tending out the end. Suppose the S_{i} are all homotopic in M \ P and π_{1}-injective in M \ P. Then e is topologically finite. ♢

Note that M is topologically finite if and only if it has finitely many ends and each end is topogically finite.

2.6. Replacing singular surfaces by embedded ones.

The following constructions are based on an argument in [FrHS]. In that paper, the results are expressed in terms of minimal surfaces, though they can also be formulated in terms of normal surface theory [JR]. In any case, the parts of the argument relevant to Theorem 2.6.2 below are purely combinatorial.

Theorem 2.6.1 : Suppose f : Σ ⟶ M is π_{1}-injective and homotopic in M to an embedding. Given any neighbourhood, U, of f(Σ) in M, there is an embedding, f′ : Σ ⟶ U, with f′ homotopic to f in M. ♢

This is a consequence of a result in [FrHS], as observed by Bonahon [Bon1] (Troisième Partie, Lemme 1.22). (We note that [FrHS] only talks explicitly about immersed surfaces, though the argument can be applied to general singular maps.) For more details, see [CanM]. ♢

Another means of obtaining embedded surfaces is via the Thurston norm. The following can be deduced using the result of Gabai [G] that the singular Thurston norm equals the non-singular norm. (See also [P].) This uses some fairly sophisticated machinery. With an additional (unnecessary) assumption on first homology, one can give a more direct argument. This is based on the proof of Theorem 2.1 of [FrHS] which is a relatively straightforward, though ingenious, tower argument. We give this as Theorem 2.6.2 below.

In what follows, we take homology coefficients in Z_{2}.

Theorem 2.6.2 : Suppose that f : Σ ⟶ M is a map inducing injections H_{1}(Σ) ⟶ H_{1}(M) and H_{2}(Σ) ⟶ H_{2}(M). Given any neighbourhood, U, of f(Σ) in M, there is an embedding S ↪ U, of another surface S, with genus(S) ≤ genus(Σ) and with H_{2}(S) ⟶ H_{2}(M) injective.

Proof : Following [FrHS] we construct a tower of double covers so that at the top of the tower we have maps h : Σ ⟶ N and λ : N ⟶ M, with f = λ ∘ h, where N is a 3-manifold that is a regular neighbourhood of h(Σ), where H_{1}(Σ) ⟶ H_{1}(N) is surjective, and where λ is locally injective. (The last statement follows from the construction, since it is assumed that h(Σ) does not lift to any double cover of N.) Moreover, we can assume that λ(N) ⊆ U. Under our hypotheses, h_{∗} : H_{1}(Σ) ⟶ H_{1}(N) is in fact an isomorphism.

Now (as with any compact 3-manifold) we have dim H_{1}(∂N) ≤ 2 dim H_{1}(N). But dim H_{1}(N) = dim H_{1}(Σ) = 2 genus(Σ), and so dim H_{1}(∂N) ≤ 4 genus(Σ). As in [FrHS], we can write ∂N = A ⊔ B, where h(Σ) homologically separates A and B in N. (In [FrHS] it was assumed that h : Σ ⟶ M is a homotopy equivalence, but only the bijectivity of h_{∗} : H_{1}(Σ) ⟶ H_{1}(N) is needed for that part of the argument. It was also assumed there that f, and hence h, is an immersion. However, the only new situation we have to consider for a general position map is the possibility that there are two paths, α, β in Σ, connecting the same pair of cross-caps in N, with h(α) = h(β). In this case, α ∪ β is null-homologous in Σ. We see that this situation is covered by the same argument as a disjoint pair of closed curves identified under h.)

Now the images of H_{2}(A), H_{2}(B) and H_{2}(S) are all equal in H_{2}(N) and hence in H_{2}(M). In particular, they are all non-zero. Also, dim H_{1}(A)+dim H_{1}(B) = dim H_{1}(∂N) ≤ 4 genus(Σ), and so we can assume that dim H_{1}(A) ≤ 2 genus(Σ). Thus, each component of A has genus at most genus(Σ). Let F ⊆ A be such a component with H_{2}(F) non-zero in H_{2}(M).

We now proceed down the tower through a sequence of double covers. If two curves, α, β ⊆ F, get identified at a given stage to a curve γ ⊆ M, then we can surger F in a neighbourhood of γ so as to eliminate the intersection and in such a way that the resulting surface is connected. Note that this preserves the class of the surface in H_{2}(M), and also preserves the Euler characteristic and hence the genus of the surface.

At the bottom of the tower, we arrive at the required embedding in U. ♢

2.7. An incompressibility condition.

Lemma 2.7.1 : Suppose that S is a compact boundary component of a 3-manifold, M, and that P ⊆ M is a closed subset. Suppose that S can be homotoped into P. Then S is π_{1}-injective in M \ P.

Proof : Passing to the cover corresponding to S, and replacing P by its preimage, we can suppose that π_{1}(M) is supported on S. We can also suppose that P is connected (taking the appropriate component). Suppose that S is compressible in M \ P. Let D ⊆ M \ P be a compressing disc for S. Let A be the component of M \ D containing P. Now π_{1}(M) splits as a free product with π_{1}(A) as a vertex group. But π_{1}(A) is all of π_{1}(M) and so this splitting must be trivial. In other words, M \ D has another component which is simply connected. Its closure in M is a 3-manifold with only 2-sphere boundary components. It follows that ∂D bounds a disc in S giving a contradiction. ♢

2.8. A covering theorem.

The following argument is due to Thurston. Another account of it can be found in [Can1].

Theorem 2.8.1 : Let N be a non-compact topologically finite aspherical atoroidal 3manifold with empty boundary. Let X be a cover of N with π_{1}(X) finitely generated. Then X is topologically finite.

Proof : Let Γ = π_{1}(N). By Thurston’s hyperbolisation theorem for Haken manifolds (see [O,Ka]), there is a convex cocompact action of Γ on H^{3} with N ≅ H^{3}/Γ. Let Y ⊆ H^{3} be the 1-neighbourhood (say) of convex hull of the limit set. Now Y/Γ is compact, and ∂Y ≠ ∅ (since N is non-compact) so there is some r ≥ 0 such that Y ⊆ N(∂Y, r). Let G = π_{1}(X) ≤ π_{1}(N) = Γ, and let Z be the 1-neighbourhood of the convex hull of the limit set of G. Thus Z ⊆ Y, and so Z ⊆ N(∂Z, r). Now ∂Z/G is homeomorphic to a component of the quotient of the discontinuity domain of G, which is compact, by Ahlfors’s Finiteness Theorem [Ah]. Thus, Z/G ⊆ N(∂Z/G, r) is compact. But int(Z/G) is homeomorphic to H^{3}/G and hence to X. ♢

3. Polyhedral constructions.

In this section, we describe the main geometric tool used in the proof. The basic idea is that if one removes certain polyhedral subsets from a hyperbolic 3-manifold the result will be negatively curved (locally CAT(−1)) in the induced path metric. In practice, these polyhedral subsets can be thought of as barriers in the sense that we can realise certain homotopy classes of maps in their complement. In this way, we construct “balanced maps” of surfaces. One can then draw various conclusions (bounded diameter lemmas etc.) much as in the case of hyperbolic 3-manifolds. Some messing around is necessary to find the appropriate spaces in which to perform these constructions to prove tameness. The details of this are discussed in Section 4. Similar ideas have been used in [Som] and in [Bow3]. (We remark that while some of the results generalise, others are specific to ambient dimension 3.) An alternative, but less intuitive, variation of these ideas is described in Section 7, where they are more readily adapted to variable curvature.

We make use of the theory of CAT(k)-spaces, see for example [BridH]. Here k will be ±1. For constant curvature, all our spaces will be locally compact. In variable curvature (Section 7) we will allow non-locally compact spaces. All the relevant facts apply equally well in that situation.

3.1. Polyhedra.

Let M be a complete hyperbolic 3-manifold. Given x ∈ M, let Δ_{x}(M) be the unit tangent space at x.

Definition : A polyhedron, Φ, in M is a locally finite embedded simplicial complex, all of whose simplices are (embedded) totally geodesic simplices.

Here, locally finite means that only finitely many simplices meet any compact subset of M, and so Φ is closed in M.

Any point x ∈ Φ determines a closed polyhedral subset, Δ_{x}(M, Φ) ⊆ Δ_{x}(M), of tangent vectors lying in Φ.

Definition : We say that Φ is balanced at x ∈ Φ, if Δ_{x}(M, Φ) is not contained in any open hemisphere of Δ_{x}(M).

Definition : We say that Φ is fat at x if Δ_{x}(M, Φ) is connected.

Definition : We say that Φ is balanced (respectively fat) if it is balanced (respectively fat) at every point.

We remark that both of these properties are closed under locally finite union.

The definition of “balanced” is essentially the same as the definition of “NLSC” given in [Can2]. It will be used here mainly to ensure that the complementary components are negatively curved. Fatness is a somewhat technical condition that ensures the complex still acts as a barrier when we take the metric completion of the complement. Without it, we could lose part of the complex altogether. For example, if Φ were a simple closed geodesic, then taking the metric completion of M \ Φ would just give us back M. We want to avoid this phenomenon.

Given a polyhedron Φ ⊆ M, write Π = Π(Φ) for the metric completion of M \ Φ, and write int Π(Φ) = M \Φ. The inclusion int Π ↪ M extends to a natural map, ω : Π ⟶ M. We write ∂Π = Π(Φ) \ Π.

If Φ is fat, then one can construct a topological (in fact, PL) collar of ∂Π in Π; that is, an arbitrarily small neighbourhood homeomorphic to ∂Π × [0, 1] with ∂Π identified with ∂Π × {0}. We see that Π(Φ) is a topological (indeed PL) manifold with boundary ∂Π. In particular, we conclude:

Lemma 3.1.1 : If Φ is fat, then int Π ↪ Π is a homotopy equivalence. ♢

We also have:

Lemma 3.1.2 : If Φ is fat and balanced, then Φ is locally CAT(−1).

Proof : This follows from the fact that for all x ∈ Φ, the completion, Δ_{x}(M, Π), of Δ_{x}(M) \ Δ_{x}(M, Φ) is globally CAT(1) (cf. [Bow3]). ♢

We have the following, slightly technical, “fattening” procedure.

Lemma 3.1.3 : Suppose that Φ is a connected compact balanced polyhedron in M, not contained in any embedded closed geodesic or geodesic segment in M. If U is any neighbourhood of Φ in M, then there is a fat balanced polyhedron, Φ′, with Φ ⊆ Φ′ ⊆ U.

Proof : If I ⊆ Φ is any maximal geodesic segment whose interior is open in Φ, we can adjoin the convex hull of V ∩ Φ, where V is a small closed polyhedral neighbourhood of I in M (interpreting in the universal cover). We apply this to all such intervals. If x ∈ Φ is any point remaining with Δ_{x}(M, Φ) not connected, then we adjoin the convex hull of V ∩ Φ for a small closed polyhedral neighbourhood, V, of x in M. ♢

For the sake of notational convenience, we will frequently assume that the map ω : Π ⟶ M is injective, and identify Π as a subset of M (namely the closure of M \ Φ). All our constructions are readily interpreted in the general case.

3.2. Polyhedral maps.

Suppose that K is a finite simplicial complex.

Definition : A map ϕ : K ⟶ M is polyhedral if the image of every simplex is a (possibly degenerate) totally geodesic simplex in M.

Here, we do not assume ϕ to be injective on simplices. However, is easily seen that the image, ϕ(K), has the structure of a polyhedron in M (possibly after subdividing, and replacing degenerate simplices by lower dimensional ones).

Definition : We say that ϕ is balanced at x ∈ K if there is some neighbourhood, V, of x in K such that ϕ(V) is a balanced polyhedron in M at ϕ(x). (Here, ϕ|V need not be injective.)

This is, of course, really a property of the induced map of the link of x into Δ_{ϕ(x)}(M).

Definition : We say that ϕ is balanced if it is balanced at every point.

We note that if ϕ is balanced, then ϕ(K) is a balanced polyhedron. Also, since it is a local property, the lift of any balanced map to any cover of M is also balanced.

We have the following simple criterion for recognising that certain maps are balanced (cf. [Bon2] Lemme 1.8). Suppose that ϕ : K ⟶ M is a polyhedral map. Suppose that x ∈ K lies in the interior of an embedded interval I ⊆ K, and that ϕ|I maps I injectively to a geodesic segment in M. Then Φ is balanced at x.

One can construct balanced maps in abundance, following the basic idea of Bonahon [Bon2] (Lemme 1.7) which elaborates an earlier idea of Thurston. We can formalise this as follows:

Lemma 3.2.1 : Suppose that K is a finite simplicial 2-complex and that we are given a homotopy class of maps from K into M. Suppose that γ ⊆ K is an embedded closed polygonal curve whose image in M is homotopically non-trivial and non-parabolic. Then (perhaps after subdividing) we can realise the homotopy class by a balanced simplicial map, ϕ : K ⟶ M, with ϕ(γ) a closed geodesic.

Proof : After subdividing, we can assume that γ lies in the 1-skeleton of K. Let ϵ be any edge of γ, and let τ be any maximal tree in the 1-skeleton of K containing γ \ ϵ. Let γ̄ be the corresponding closed geodesic in M, and choose any y ∈ γ̄. (In general, we should allow γ̄ to wrap a number of times around a primitive geodesic in M.) We now fix a relative homotopy class (K, τ) ⟶ (M, y) in the right free homotopy class, such that (τ ∪ ϵ, τ) gets sent to the class (γ̄, y). (There is some choice in “spinning” around γ̄, but this need not concern us here.) We now define ϕ by sending τ to y, and sending each remaining edge of the 1-skeleton to the corresponding geodesic loop based at y (so that ϕ(ϵ) = γ̄). We send each 2-simplex to a geodesic simplex in M. ♢

Of course, ϕ(K) will not be in any sense convex, though we can include additional geodesic loops into the construction by the following procedure (the eventual aim of which is to exhaust the convex core up to bounded distance, see Section 5).

Suppose that K is a 2-complex and α is a polygonal path in K (not necessarily embedded). After subdivison of K, we can suppose that α lies in the 1-skeleton. We form another complex K_{α} by gluing a disc to α along a subarc of its boundary. We write α̂ for the unattached part of the boundary of the disc. Thus, we can think of the disc as realising a homotopy from α̂ to α in K_{α}, relative to their common endpoints. We refer to the disc we have added as a “fin” attached to K.

Lemma 3.2.2 : Suppose that ϕ : K ⟶ M be a balanced polyhedral map, and let α be a path in K. We can extend ϕ to a balanced simplicial map ϕ : K_{α} ⟶ M, so that ϕ(α̂) is geodesic in M.

Proof : We can realise our fin by taking triangulation whose 1-skeleton is a zigzag between the paths α and α̂ (cf. the construction of [Bon2] Section II). ♢

3.3. Polyhedral maps into Π(Φ).

In the above, we assumed the codomain of our maps to be M. However many of the essential principles apply when this is replaced by a locally finite locally CAT(−1) polyhedral complex. Here we restrict our attention to spaces of the form Π(Φ) as described above. In order to quote results from elsewhere, we note that such spaces can be triangulated as polyhedral complexes, though this construction is a little artificial, given that the essential arguments involving such complexes adapt directly to this set up.

For simplicity of exposition, we shall assume Π ⊆ M to be embedded, so that for all y ∈ Π, we have Δ_{y}(M, Π) ⊆ Δ_{y}(M). Since Φ is balanced, Δ_{y}(M, Π) is globally CAT(−1) in the induced path metric.

We need to define “balanced” maps in this context. One can give a similar definition as before, though we shall restrict here to the case where the domain is a circle or a surface. Let us suppose first that our maps are non-degenerate, i.e. each simplex maps locally injectively.

Suppose that α is a circle, and that ϕ : α ⟶ Π is a (non-degenerate) polyhedral map. Each point x ∈ α determines a pair of points in Δ_{ϕ(x)}(M, Π). The condition that α is locally geodesic at x is equivalent to asserting that these points are distance at least π apart in the induced path metric on Δ_{ϕ(x)}(M, Π). In this case we deem ϕ to be balanced at x.

Suppose now that Σ is a closed surface, and that ϕ : Σ ⟶ Π is a polyhedral map. This time, each x ∈ Σ determines a closed path in Δ_{ϕ(x)}(M, Π). For us, the key point is that this curve should have length at least 2π. This is satisfied, for example, if it contains two points distance at least π apart in the induced path metric on Δ_{ϕ(x)}(M, Π). This is in turn satisfied if x is contained in the interior of an interval Σ that gets mapped injectively to a local geodesic segment in Π. Such a map will deemed to be “balanced”.

If ϕ maps each simplex to a simplex of the same dimension (so that after subdividing, ϕ can be assumed injective on each simplex) then the pull-back metric on Σ is hyperbolic with cone singularities at least 2π. This fact will be used to bound the area of Σ, via the Gauss-Bonnet theorem (see Section 3.4).

In general, we may be faced with the prospect that some simplices will collapse to lower dimensional ones. In this case, we need to consider neighbourhoods of subcomplexes that get collapsed to points. Essentially the same reasoning goes through. In this case we get a pseudometric on Σ whose hausdorffification is singular hyperbolic. Again we will need a bound on the area. Perhaps the simplest way to deal with the technicalities is to view the Gauss-Bonnet theorem in this context as essentially combinatorial: summing angles of each triangle and using Euler’s formula directly.

We can construct polyhedral maps of this type exactly as with Bonahon [Bon2] (cf. Lemma 3.2.1 above). To extend over 2-simplices, some care is needed. We can construct a ruled surface by coning an edge over an opposite vertex (modulo some technicalities, this would suffice for our purposes). If we want a bona fide polyhedral map, then we may need to subdivide into smaller simplices.

We remark that the composition of a balanced map Σ ⟶ Π with the inclusion of Π into M need not be balanced as a map into M. (One can construct examples where the closed path in Δ_{ϕ(x)}(M, Π) is contained in an open hemisphere in Δ_{ϕ(x)}(M).) However, this need not matter to us, other than as a point of caution.

3.4. The thick-thin decomposition.

First, we recall the thick-thin decomposition of a complete hyperbolic 3-manifold M without parabolics.

Let η > 0 be some fixed constant less than the Margulis constant. Let T (M) be the set of x ∈ M such that x lies in some essential curve β_{x} in M of length at most η. We refer to T (M) as the thin part of M. Given a primitive free homotopy class, α, of closed curve in M, let T (M, α) be the set of x for which β_{x} can be chosen to be homotopic to α^{n} for some n ∈ Z \ {0}. Such a “Margulis tube” is a uniform neighbourhood of the corresponding closed geodesic, α_{M}, in M. The thin part, T (M), is a disjoint union of T (M, α) as α varies over such primitive classes. The write Θ(M) for the thick part of M, that is, the closure of M \ T (M). Let ρ be the riemannian path pseudometric that agrees locally with the hyperbolic metric, d, on Θ(M), and which is set to be zero on each Margulis tube, often termed the “electric pseudometric”.

The following is a generalisation of the well known “bounded diameter lemma” where we allow for surfaces that are not π_{1}-injective.

Lemma 3.4.1 : Suppose that f : (Σ, σ) ⟶ (M, d) is a 1-lipschitz map from a closed hyperbolic surface of curvature at most −1 to M. There is a constant, k, depending only on genus(Σ) and η, such that either the ρ-diameter of f(Σ) is less than k, or else, for any x ∈ f(Σ), there is a homotopically non-trivial simple closed curve, β, of length at most η in Σ, with f(β) homotopically trivial in M and with ρ(x, f(β)) < k.

Proof : This follows by a standard argument. Briefly, each component of the thick part of Σ has bounded σ-diameter. We can pass between any two points in Σ passing through a bounded number of components of the thin part of Σ. Each component of the thin part either maps into a Margulis tube of M and is therefore not noticed by the pseudometric ρ, or else is homotopically trivial in M, giving rise to our curve β. ♢

4. The wrapping construction.

This section contains the core of the argument. We give our reinterpretation of the shrinkwrapping construction (cf. [CalG,Som]), to equip us with suitable bounded genus surfaces in a 3-manifold. How this construction is used will be explained in Section 5.

Let M be a complete hyperbolic 3-manifold without cusps, and suppose that e is an end of M carrying the fundamental group of M (as in Section 2.4). In other words, π_{1}(M), is a compression group, and any Scott core will be a compression body, with e as the outer end. We write e_{1}, …, e_{n} for the (possibly empty) set of inner ends, each of which is incompressible. If e itself is not incompressible, we shall assume that each of the inner ends is topologically finite.

Let (g_{0}, g_{1}, …, g_{n}) be the type of e. Thus, for i ≠ 0, g_{i} = genus(e_{i}). We set g = genus(e) = ∑_{i=0}^{n} g_{i}.

Let K be a finite simplicial complex of the homotopy type of M. (For example, take a wedge of n surfaces Σ_{i} of genus g_{i}, together with g_{0} loops.) By Lemma 3.2.1, there is a balanced polyhedral homotopy equivalence, ϕ : K ⟶ M.

Suppose that α is a path in K, and let K_{α} be the complex constructed in Section 3.2. We can extend ϕ to a balanced homotopy equivalence K_{α} ⟶ M, also denoted ϕ, which sends α̂ to a geodesic path.

We will also take as hyphothesis the existence of a proper ray, τ, based in ϕ(K_{α}), and with ρ(τ, ϕ(K)) sufficiently large, to be specified later. Here ρ is the electric pseudometric on M as defined in Section 3.

We can now fatten ϕ(K) and ϕ(K_{α}) to give fat balanced polyhedral complexes Φ ⊆ Φ_{α} ⊆ M (Lemma 3.1.3). To save overburdening our notation, we shall assume that these complexes are already fat, and write Φ = ϕ(K) and Φ_{α} = ϕ(K_{α}). This does not affect our argument. Again, for notational convenience, we regard Π(Φ) and Π(Φ_{α}) as subsets of M.

Since H_{2}(K) ⟶ H_{2}(M) surjects, Φ separates the ends of M. The same is true of Φ_{α}. Let Υ and Υ_{α} be the components of Π(Φ) and Π(Φ_{α}) respectively containing the end e. Since H_{1}(Φ) ⟶ H_{1}(M) also surjects, its boundary ∂Υ = Υ ∩ Φ is connected. Now Υ is a full neighbourhood of this end, so that the discussion of Section 2.4 applies. This is also true of Υ_{α}.

The main aim of this section is to prove the following statement:

Proposition 4.1 : Suppose ϕ : K_{α} ⟶ M and K ⊆ K_{α} are as above. Suppose that τ is a proper ray going out the end e, with basepoint in K_{α}, and with ρ(τ, ϕ(K)) sufficiently large (depending on the geometry of Υ). Then there is an embedded surface S ⊆ int Υ, with genus(S) = g, with S incompressible in Υ, and separating the end e from ∂Υ, and such that both ρ(τ, S) and the ρ-diameter of S are bounded in terms of g and the 3-dimensional Margulis constant.

We shall explain in Section 5 how this is used to give tameness of e. We now set about the proof of Propostion 4.1.

Let F ⊆ int Υ_{α} be an embedded surface separating Φ_{α} from the end e and with π_{1}(F) ⟶ π_{1}(Υ_{α}) injective. Note that such a surface exists — take any surface that separates of minimal possible genus, and use Dehn’s lemma and the fact that ∂Υ_{α} is connected to see that it must be incompressible in Υ_{α}. We shall take F to be smooth.

From this point on, we split the discussion into two cases. First we consider the case where e is not incompressible, and that each of the ends e_{i} is topologically finite. We will describe the modifications necessary for the incompressible case later.

Let W be the component of M \ F containing Φ_{α}. Note that each of the inner ends, e_{i}, of M is also an end of W. Thus W is topologically finite. Clearly W is aspherical (since M is). In fact:

Lemma 4.2 : W is atoroidal.

In fact, this lemma can be bypassed, but at the cost of making appeal to some more sophisticated results, as is done in [Som]. In order not to detract from the main flow of the argument, we postpone the proof until the end of this section.

Recall that we have maps K ↪ K_{α} ⟶^{ϕ} W ↪ M and that the composition K ⟶ M induces an isomorphism on π_{1}. It follows that π_{1}(K) ⟶ π_{1}(W) and π_{1}(K_{α}) ⟶ π_{1}(W) inject. They have the same image G ≤ π_{1}(W). Let X be the cover of W corresponding to G. Thus the map X ⟶ M obtained by composing X ⟶ W ⟶ M is a homotopy equivalence.

Lemma 4.3 : X is topologically finite.

Proof : Since G ≅ π_{1}(X) is finitely generated, this follows from Lemma 4.2 and Theorem 2.8.1. ♢

Since the inner ends of W correspond to surface group factors of G, they lift to ends of X, also denoted by e_{i}.

We will later need the following observation:

Lemma 4.4 : Suppose that α, β are primative essential closed curves in X which project to the same closed curve in W. Then α = β.

Proof : Since the composition X ⟶ W ⟶ M is a homotopy equivalence, α and β are freely homotopic in X. The annulus realising this homotopy projects to a (singular) torus in W. Since W is atoroidal (Lemma 4.2), this torus bounds a (singular) solid torus in W with core curve homotopic to the common image of α and β (since these curves are assumed primitive and non-trivial). Thus image of the annulus in W can be homotoped relative to its boundary to the curve that is common image of α and β. Since X ⟶ W is a covering space, this homotopy lifts to X showing that α = β. ♢

(Again, Lemma 4.2 could be bypassed using the observation that α and β are nontrivial in X and hence in M.)

Let ϕ̃: K_{α} ⟶ X be the lift of ϕ : K_{α} ⟶ W. Thus, ϕ̃ and of ϕ̃|K are both homotopy equivalences. We write Φ̃ ⊆ Φ̃_{α} respectively for the fattenings of ϕ̃(K_{α}) and of ϕ̃(K). For simplicity we will pretend that these are already fat. (This avoids some technical fussing, but makes no essential difference to the argument.) We write Ψ ⊆ Ψ_{α} for the preimages of Φ and Φ_{α} under the covering map X ⟶ W. Thus Φ̃ ⊆ Ψ and Φ̃_{α} ⊆ Ψ_{α}. All these polyhedra are fat and balanced in X. (These definitions are local, and so make sense also in an incomplete hyperbolic manifold such as X.) Note that Φ̃ and Φ̃_{α} are compact.

Since K_{α} ⟶ X is a homotopy equivalence, π_{1}(X) is a compression group. Now X is topologically finite, and we have observed that each surface factor of G ≅ π_{1}(X) is conjugate to an end of X. Thus, by Lemma 2.2.1, X must be homeomorphic to the interior of a compression body of type (g_{0}, g_{1}, … g_{n}). Its inner ends correspond precisely to the inner ends, e_{1}, …, e_{n}, of M. We can find a manifold P ⊆ X, containing Φ̃_{α} = f̃(K_{α}), with X \ P homeomorphic to ∂P × [0, ∞). (Imagine pushing the outer boundary of the manifold compactification of X slightly inside X.) Thus, ∂P is a surface of genus g. Note that ϕ̃: K_{α} ⟶ P is again a homotopy equivalence.

While X is topologically nice, its intrinsic geometry may be complicated, since we have not made any geometric assumptions about the boundary, ∂W = F, of W, and this issue will arise in the cover, X. For this reason, our next job is will be to embed X in a locally CAT(−1) polyhedral complex Z. In fact, Z will be a hyperbolic 3-manifold with polyhedral boundary, ∂Z. We will carry out the “wrapping” procedure in Z. (The idea is based on a similar construction in [Som].)

Recall that F = ∂W is embedded incompressible surface in Υ_{α}, where Υ_{α} is the component of Π(Φ_{α}) containing the end e. We are assuming that ∂W is smooth (largely for convenience of terminology). Note that W ∪ ∂W is the metric completion of W. We write X ∪ ∂X for the metric completion of X. This is a manifold with smooth boundary ∂X, and the covering map X ⟶ W extends to a covering map X ∪ ∂X ⟶ W ∪ ∂W.

Let Υ̃_{α} be the universal cover of Υ_{α}. This has polyhedral boundary, ∂Υ̃_{α}. Let F̃ be a component of the lift of F. Since F ↪ Υ_{α} is π_{1}-injective, F̃ is a universal cover of F. Now, F is 2-sided, with “outside” on the side of e. Let A be the component of Υ̃_{α} \ F̃ on the outside of F̃. By van-Kampen, this is also simply connected. Note that it is also aspherical (since Υ_{α} is). The group π_{1}(F) acts on A.

Suppose that F̂ is a any covering space of F. We get a quotient, AF̂, of A by the action of π_{1}(F̂). We can identify the quotient of F̃ with F̂. Thus AF̂ is a 3-manifold with smooth boundary, F̂, as well as a polyhedral boundary (namely the quotient of A ∩ ∂Υ̃_{α}). Also AF̂ is aspherical, so the inclusion F̂ ↪ AF̂ is a homotopy equivalence. Thus AF̂ deformation retracts onto F̂.

Now ∂X ⟶ ∂W = F is a union of covering spaces. If F̂ is a component of ∂X, we can glue a copy of AF̂ to X along F̂. Since X lies on the “inside” of F̂, we get a seamless gluing. In other words, X ∪ AF̂ is locally isometric to H^{3} in a neighbourhood of F̂. Since AF̂ is a subset of a covering of Υ_{α}, there is a natural map λ : Â ⟶ Υ_{α} ⊆ M. Note that λ(int Â) ∩ Φ_{α} = ∅.

Performing this construction for each component of ∂_{0}X, we get our locally CAT(−1) space, Z. Note that ∂X is precisely the topological boundary of X in Z, and that Z deformation retracts onto X. Combining the various maps λ with the covering X ⟶ W ⊆ M, we get a natural map Z ⟶ M, also denoted λ. By construction, λ is 1-lipschitz. Writing int Z = Z \ ∂Z for hyperbolic part of Z, we have λ^{−1}(Φ) ∩ int Z = Ψ ⊆ X and λ^{−1}(Φ_{α}) ∩ int Z = Ψ_{α} ⊆ X. Since Φ̃ ⊆ W maps surjectively to Φ ⊆ W ⊆ M, we have λ(Φ̃) = Φ. Similarly, λ(Φ̃_{α}) = Φ_{α}. Note that the inner ends of X can be identified with the inner ends of M, and are not affected by the construction of Z. In particular, λ is injective on the inner ends of Z.

In summary, we have the following spaces and maps:

P ↪ X ↪ Z
↗  ↓  ↓
K ↪ K_{α}  →  W ↪ M

The maps K ↪ K_{α} ⟶ P ↪ X ↪ Z ⟶ M are all homotopy equivalences. The space Z is locally CAT(−1), and λ|X : X ⟶ W is a covering space. Also, Z deformation retracts onto X and X onto P. We have λ^{−1}(Φ_{α})∩int Z = Ψ_{α} ⊆ X. The map int Z\Ψ_{α} ⟶ M \Φ_{α} is a covering space (with possibly disconnected domain). The map λ : Z ⟶ M is 1lipschitz.

The plan now is to homotope ∂P in the completion of Z \ Φ̃_{α} to a balanced map, and then project down to M. As observed in Section 1, this homotopy may sweep through Φ, so we need to check that the resulting surface separates Φ from the end e in M.

For this purpose, we bring our ray, τ, into play. Recall that τ ⊆ M with basepoint in Φ_{α} and goes out the end e. Moreover, we are assuming that ρ(τ, Φ) is sufficiently large (to be specified shortly). We can assume that τ meets Φ_{α} only in its basepoint, and so lifts to a proper ray τ̃ in Z with λ(τ̃) = τ. (Recall that λ restricted to the preimage of M \ Φ_{α} is a covering space.) Now τ̃ is a proper ray in Z that eventually leaves X, and hence P.

Thus it crosses ∂P essentially. (More formally, the locally finite Z_{2}-homology cycle given by τ̃ has non-trivial intersection with the class of ∂P in H_{2}(Z).)

Now Φ̃_{α} is fat and balanced in Z, and so we can form the space Π_{Z}(Φ̃_{α}) — the completion of Z \ Φ̃_{α} — exactly as for manifolds (see Section 3.2). For notational convenience, we shall assume, as usual, that Π_{Z}(Φ̃_{α}) is a subset of Z. We write Y_{α} for the component containing ∂P. Thus, Y_{α} is a hyperbolic 3-manifold with polygonal boundary ∂Y_{α}. Note that τ̃ ⊆ Y_{α}. We can similarly define Y as a component of Π_{Z}(Φ̃). Note that, since Φ̃ carries all of H_{1}(Z) and H_{2}(Z), we see from the discussion of Section 2.3 that the topological boundary of Y in Z is connected. This topological boundary is just Y ∩ Φ̃. Moreover, Y minus its other manifold boundary components has exactly one end, separated from Φ̃ ∩ Y by ∂P.

By construction, ∂P is homotopic in Z into Φ̃_{α}. Thus, by Lemma 2.7.1, it is incompressible in Z \ Φ̃_{α} and hence also in Υ_{α}. Note that ∂P is also incompressible in Z \ Φ̃ and hence Y for the same reason.

We can now find a balanced map h̃ : Σ ⟶ Y_{α}, homotopic in Y_{α} to ∂P, where Σ is a surface of genus g. The induced pseudometric, σ, on Σ is singular hyperbolic. The composition h = λ ∘ h̃ is 1-lipschitz. Note that, h̃(Σ) must also intersect τ̃ homologically in Υ̃_{α}. In particular, τ̃ ∩ h̃(Σ) ≠ ∅, and so τ ∩ h(Σ) ≠ ∅.

Lemma 4.5 : Provided ρ(τ, Φ) is big enough, we have h(Σ) ∩ Φ = ∅, and the ρ-diameter of h(Σ) is bounded (depending only on genus(Σ) and the Margulis constant).

Proof : Write η_{0} for the 3-dimensional Margulis constant. By Lemma 3.4.1, the only way this might fail would be if there were an essential curve, β ⊆ Σ, of σ-length at most η_{0}, with h(β) homotopically trivial in M, and with ρ(τ, h(β)) bounded by some constant k (depending on genus(Σ) and η_{0}). Now λ : Z ⟶ M is a homotopy equivalence, and so h̃(β) is homotopically trivial in Z. But Z is locally CAT(−1), and so (coning over some point of h̃(β)), it bounds a (singular) disc, D, in Z, of diameter at most η_{0}. Thus λ(D) is a disc of diameter at most η_{0} bounding h(β) in M. Now, provided that ρ(τ, Φ) > k + 2η_{0}, then ρ(h(β), Φ) > η_{0} and so λ(D) ∩ Φ = ∅. Since λ(Φ̃) = Φ, it follows that D ∩ Φ̃ = ∅, and so h̃(β) is homotopically trivial in Z \ Φ̃. But h̃(Σ) is homotopic to ∂P in Π(Φ̃_{α}) and so, in particular, in Π(Φ̃). Thus, h̃ : Σ ⟶ Z \ Φ is π_{1}-injective, giving a contradiction. ♢

In other words, h(Σ) ⊆ int Υ. It is still not clear that h(Σ) is homologically non-trivial in Υ. To make life simpler, we first replace it with another surface, f(Σ), as follows.

By [FrHS] (Theorem 2.6.1 here), h̃ is homotopic in Y to an embedding f̃: Σ ⟶ int Y with f̃(Σ) in an arbitrarily small neighbourhood of h̃(Σ). Let f = λ ∘ f̃: Σ ⟶ M. From Lemma 4.5, we see that f(Σ) ∩ Φ = ∅, and f(Σ) has bounded ρ-diameter.

Let Q be the component of Z \ f̃(Σ) containing Φ̃ = ϕ̃(K). The topological boundary ∂Q = f̃(Σ) is homotopic to ∂P in Y. Now, as observed above, Y has Φ̃∩Y as one boundary component. Writing Y′ for Y minus the other (polyhedral) boundary components, we see that Y′ has one end, separated from Φ ∩ Y by the surface ∂P. Considering their homology classes in H_{2}(Y), it follows that ĥ(Σ) and hence also ∂Q = f̂(Σ) separate Φ ∩ Y from the end. Since Σ has genus g — the same genus as the end, it follows by Lemma 2.4.3, that the end of Y′ deformation retracts onto ∂Q. Since K ⟶ int Z is a homotopy equivalence, and factors through ϕ̃: K ⟶ Q, it follows K ⟶ Q ⟶ int Z are also homotopy equivalences. Thus Q ⟶ Z is a homotopy equivalence. In particular, π_{1}(Q) is a compression group, and so Q in the interior of a compression body with outer boundary ∂Q, and inner ends e_{1}, …, e_{n}.

Recall that Ψ = λ^{−1}(Φ) ∩ int Z. Since λ is injective on the inner ends of Z, we see that Ψ ∩ Q is compact. We claim:

Lemma 4.6 : Ψ ∩ Q = Φ̃.

Proof : By construction, Φ̃ ⊆ Ψ ∩ Q. Note that Ψ ∩ ∂Q = ∅, and so Ψ ∩ Q is a union of compact components of Ψ. Now Ψ is the preimage of Φ = ϕ(K) under the covering map X ⟶ W ⊆ M. Thus Ψ ∩ Q is a finite union, Φ̃(K) ∪ ⋃_{i=1}^{q} ϕ̃_{i}(K_{i}), where each K_{i} is a finite cover of K, and where λ ∘ ϕ̃_{i} is equal to the covering K_{i} ⟶ K postcomposed with ϕ.

Suppose x ∈ K_{i}. We can assume that π_{1}(K) is non-cyclic, and so it’s not hard to find a loop β in K_{i} based at x which maps to a non-trivial primative loop in K under the finite covering K_{i} ⟶ K (for example, using the fact that π_{1}(K) contains a free group of rank 2). Now λ(ϕ̃_{i}(β)) = λ(ϕ̃(γ)), and so, by Lemma 4.4, we have ϕ̃_{i}(β) = ϕ̃(γ). In particular, ϕ̃_{i}(x) ∈ ϕ̃(K). Thus, ϕ̃_{i}(K_{i}) ⊆ Φ̃. This shows that Ψ ∩ Q = Φ̃ as claimed. (In fact, one can see there can be no such ϕ̃_{i}.) ♢

In other words, we have shown that Φ̃ is the entire preimage of Φ = ϕ(K) under the map λ|Q, so that λ|Φ̃ : Φ̃ ⟶ Φ is a homeomorphism.

In what follows, we take homology with Z_{2} coefficients.

Lemma 4.7 : The map f : Σ ⟶ M \ Φ induces an injection H_{2}(Σ) ⟶ H_{2}(M \ Φ).

Proof : In other words, we claim f(Σ) is non-trivial in H_{2}(M \ Φ). We distinguish two cases.

The easy case is where Q is not a handlebody. Then f̃(Σ) = ∂Q is non-trivial in H_{2}(Q). Since λ : Q ⟶ M is a homotopy equivalence, f(Σ) = λ(f̃(Σ)) is also non-trivial in H_{2}(M) and so, in particular, in H_{2}(M \ Φ).

Suppose that Q is a handlebody. Choose any point x ∈ Φ̃ ⊆ Q. Then ∂Q = f̃(Σ) is homologically linked with x in Q. Now (λ|Q)^{−1}(λ(x)) = {x}. Since ∂Q is compact, and λ is locally injective, it follows that λ(∂Q) = f(Σ) is homologically linked with λ(x) in M. (The 3-chain λ(Q) has boundary λ(∂Q) and has non-zero Z_{2}-intersection with λ(x).) It follows that λ(∂Q) = f(Σ) is non-trivial in H_{2}(M \ {λ(x)}) and hence also in H_{2}(M \ Φ), since λ(x) ∈ λ(Φ̃) = Φ. ♢

Lemma 4.8 : The map f : Σ ⟶ M \ Φ induces an injection H_{1}(Σ) ⟶ H_{1}(M \ Φ).

Proof : Since ∂Q = f(Σ) and f = λ ∘ f̃, this is the same as saying that the map λ : ∂Q ⟶ M \ Φ induces an injection to H_{1}(M \ Φ).

Let a ∈ H_{1}(∂Q) with λ(a) trivial in H_{1}(M \ Φ). Let b be a 2-chain in M \ Φ with ∂b = λ(a).

Now since λ : Q ⟶ M is a homotopy equivalence, and λ(a) is trivial in M, a must be trivial in H_{1}(Q). It thus bounds a non-trivial relative 2-cycle, c, in H_{2}(Q, ∂Q). By Alexander duality, there is some d ∈ H_{1}(Q) with Z_{2}-intersection ⟨c, d⟩ = 1.

Now the 2-chains b and λ(c) both have the same boundary, λ(a), in M, and so e = b ∪ λ(c) gives us a cycle representing an element of H_{2}(M).

Recall that ϕ̃ : K ⟶ Q is a homotopy equivalence, and so d = ϕ̃(h) for some h ∈ H_{1}(K). Thus λ(d) = λ(ϕ̃(h)) = ϕ(h), which is supported on ϕ(K) = Φ. Since b does not meet Φ, and since (λ|Q)^{−1}Φ = Φ̃ = ϕ̃(K) and λ|Φ̃(K) is injective, we see that ⟨e, λ(d)⟩ = ⟨b ∪ λ(c), λ(d)⟩ = ⟨λ(c), λ(d)⟩ = ⟨c, d⟩ = 1 where the first three angle brackets ⟨, ⟩ denote the intersection number on H_{2}(M) × H_{1}(M).

But H_{2}(M) is generated by the homology classes h(e_{1}), …, h(e_{n}) of its inner ends e_{1}, …, e_{n}, and so the intersection form on H_{2}(M)×H_{1}(M) is trivial, giving a contradiction.

♢

We have shown that f : Σ ⟶ M\Φ is injective on both H_{1} and H_{2}. We can thus apply Lemma 2.6.2 to give us an embedded surface S in M \ Φ, with genus(S) ≤ genus(Σ) = g, with S in a small neighbourhood of f(Σ) and non trivial in H_{2}(M \ Φ). (We remark that we could bypass Lemma 4.8, using the Thurston norm [G,P].) In particular, ρ(τ, S) and the ρ-diameter of S are both uniformly bounded. Note that S is contained in the component, Υ, of Π(Φ) containing e. Since ∂Υ = Φ ∩ Π(Φ) is connected, and e is the only end of Υ, it follows that S must separate e from ∂Υ. Since genus(e) = g = genus(Σ), we have genus(S) = g by Lemma 2.4.4.

This proves Proposition 4.1 in the compressible case.

The incompressible case calls for a few minor modifications. In this case, M, has two ends e and e′. We cannot assume that e′ is topologically finite. Now Φ_{α} separates e and e′, and in this case, we take incompressible surfaces, F and F′, in M \ Φ_{α} separating Φ_{α} from e and e′ respectively. Let W ⊇ Φ_{α} be the relatively compact open submanifold of M with ∂W = F ⊔ F′. In this case also, W is atoroidal (by essentially the same argument) and is therefore covered by a topologically finite manifold X. We find a compact manifold P ⊆ X, with Φ_{α} ⊆ P ⊆ X and with X \ P just a product. In this case P is homeomorphic to a product Σ × [0, 1], where Σ is a surface of genus g. (This follows from the fact that a manifold with fundamental group a surface group is just a product.) Now τ crosses one of the boundary components, ∂_{0}P, of ∂P essentially. We proceed as before, with ∂_{0}P playing the role of ∂P. In this case, ∂_{0}P ⟶ M is a homotopy equivalence, and so most of the topological reasoning is simpler. In particular, Lemmas 4.7 and 4.8 are immediate.

This proves Proposition 4.1 in the incompressible case.

We now return to:

Proof of Lemma 4.2.

We want to prove that W is atoroidal. An alternative argument in a slightly different set-up is given in [CalG].

We start with some general lemmas.

Suppose that Ω ⊆ H^{3} is a (perhaps disconnected) locally finite balanced polyhedron.

Lemma 4.9 : Suppose V ⊆ H^{3} is a (connected) codimension-0 submanifold with Ω ∩ ∂V = ∅ and with ∂V incompressible in H^{3} \ Ω. Then π_{1}(V) is free.

To begin the proof, we fix a basepoint, x, and set d_{x}(y) = d(x, y). We can assume that all surfaces in H^{3} are in general position with respect to d_{x}, i.e. the restriction of d_{x} to the surface is a Morse function. Our aim is to show that we can isotope ∂V in H^{3} \ Ω so that d_{x}|∂V has no local maximum. We use:

Lemma 4.10 : Suppose that t > 0 and that F ⊆ H^{3} \ Ω is a compact surface with ∂F ⊆ ∂N(x, t), and which is π_{1}-injective in H^{3}. Suppose that F_{0} ⊆ F \ int N(x, t) is connected. Then there is a closed surface J ⊆ F containing F_{0}, with ∂J ⊆ ∂N(x, t) which is isotopic in H^{3} \ Ω relative to ∂J to a compact surface J′ ⊆ ∂N(x, t) \ Ω.

Proof : Let us first consider the case where F = F_{0}. Write R = H^{3} \ (Ω ∪ int N(x, t)) and given u > t, write R(u) = R ∩ int N(x, u). These are manifolds with boundary ∂R = ∂N(x, t) \ Ω. There is some u_{0} > t such that F ⊆ R_{0} = R(u_{0}).

Since Ω is balanced, the function d_{x} has no local maximum on Ω. From this, it follows that, as u increases from t to u_{0}, the topology of R(u) changes a finite number of times by the addition of 0-handles or 1-handles (never 2-handles). Thus R_{0} is homeomorphic to ∂R_{0} × [0, 1) with a finite number of (open) 1-handles attached. In particular, it follows that H_{2}(R_{0}, ∂R_{0}) = 0, and that π_{1}(R_{0}) is supported on the end of R_{0}. (In other words, if H ⊆ R_{0} is compact, then π_{1}(R_{0} \ H) ⟶ π_{1}(R_{0}) surjects.)

From the first observation, we see that F cuts off a compact manifold H ⊆ R_{0}. By assumption, π_{1}(F) ⟶ π_{1}(R_{0}) injects, so we can write π_{1}(R_{0}) = π_{1}(R_{0} \ H) ∗_{π_{1}(F)} π_{1}(H). But π_{1}(R_{0} \ H) ⟶ π_{1}(R) surjects, and so the splitting is trivial, i.e. π_{1}(F) ⟶ π_{1}(H) surjects. Thus F ↪ H induces a bijection on π_{1}, and so (H, F) ≅ (F × [0, 1], F × {0}) by (see [H]). Thus, F is isotopic to a surface J′ ⊆ ∂R as required.

Now consider the general case. If F_{0} were incompressible in R, then we could apply the above observation with J = F_{0}. If not, let D_{1} ⊆ R be a compressing disc for F_{0} in R. Since F is incompressible in H^{3} \ Ω, ∂D_{1} bounds a disc D′_{1} ⊆ F. Now the 2-sphere D_{1} ∪ D′_{1} bounds a ball in H^{3} which does not meet Ω. This means we can isotope D′_{1} to D_{1} in H^{3} \ Ω fixing ∂D_{1}. Now surgering F_{0} along D_{1} and continuing in this way, we get a sequence of compressing discs D_{1}, …, D_{p}, so that the surface J = F_{0} ∪ D′_{1} ∪ ⋯ ∪ D′_{p} ⊆ F is isotopic to an incompressible surface in R. By the first paragraph, this is now isotopic to a compact surface J′ ⊆ ∂R = ∂N(x, t) \ Ω, as required. ♢

Lemma 4.11 : Suppose that S ⊆ H^{3} \ Ω is a properly embedded π_{1}-injective surface. Then we can isotope S in H^{3} \ Ω so that d_{x}|S has no local maximum.

Proof : This is the same as requiring that no component of S \ N(x, t) is relatively compact for any t > 0.

If not, let t > 0 be minimal so that S \ N(x, t) has a relatively compact component. Its closure, F_{0}, satisfies ∂F_{0} = F_{0} ∩ N(x, t) = F_{0} ∩ ∂N(x, t) and meets the closure of the remainder of S\int N(x, t) at a single point, z ∈ ∂F_{0} ⊆ ∂N(x, t). Let F ⊆ F_{0} be the surface with any disc components of S \ F adjoined. By Lemma 3.10, we can find some surface J ⊆ F, with J ⊇ F_{0}, which is isotopic to some J′ ⊆ ∂N(x, t) \ Ω relative to ∂J. Since t is minimal, J′ ∩ S = ∅ (otherwise J′ would cut off a compact component of S \ int N(x, t), sandwiched between J and J′, contradicting the minimality of t). Now z ∈ ∂J′. and we can tilt J′ slightly so that d_{x}|J′ has a unique maximum at z. By replacing J by this surface, we can eliminate the critical point z. We then proceed to the next smallest value of t for which S \ N(x, t) has a relatively compact component. Since this set is discrete, we eventually isotope S so that d_{x}|S has no maximum, as required. ♢

Proof of Lemma 4.9 : By Lemma 4.11, we can isotope V so that d_{x}|∂V has no local maximum. Now as t increases, the topology of the manifold V ∩ N(x, t) can change only by the addition of 0-handles or 1-handles. The result follows. ♢

We can now proceed with the proof of Lemma 4.2.

Suppose, for contradiction, that Δ is an embedded essential torus in W. Now Δ cannot be essential in M, and so it has a compressing disc in M. Since M is aspherical, it follows that Δ bounds either a solid torus or a ball with knotted hole in M. In the former case, the solid torus would have to contain F = ∂W, which easily leads to a contradiction. We can therefore assume that Δ bounds a ball with knotted hole. In particular, Δ is homotopically trivial in M and so lifts to H^{3}. Let V be the component of the preimage of W in H^{3} that contains Δ. Note that Δ is also incompressible in V. Let Ω be the preimage of Φ_{α} in H^{3}. Note that ∂V is incompressible in H^{3} \ Ω. By Lemma 4.9, π_{1}(V) is free. But π_{1}(Δ) ⟶ π_{1}(V) is injective, giving a contradiction.

This completes the proof.

5. Proof of tameness when there are no cusps.

Let M be a complete hyperbolic 3-manifold without cusps and with π_{1}(M) finitely generated. Let e be an end of M. We show:

Proposition 5.1 : e is topologically finite.

We write core(M) for the convex core of M.

Definition : We say that e is geometrically finite if there is a neighbourhood, E, of e with E ∩ core(M) = ∅. Otherwise, we say that e is degenerate.

If e is geometrically finite, then we can take E to be bounded by a convex surface that is a boundary component of core(M). In this case, topological finiteness follows easily.

The real interest is in the degenerate case. In other words, we have a sequence x_{i} ∈ core(M) with x_{i} → e. (We remark that, using Ahlfors’s finiteness theorem, one can see, in fact, that core(M) contains a neighbourhood of the end. However, we shall not use this, since we want an argument that is adaptable to the case of variable curvature where this is not known a-priori.)

For the purposes of the proof, we split the degenerate case into two subcases, depending on whether e is incompressible or compressible. We shall assume the incompressible case in dealing with the compressible case.

Before continuing, we make the following observation (cf. Section 1.5 of [Bon2]).

Lemma 5.2 : There is some constant k_{0} such that for any b ∈ M, any point of core(M) lies a distance k_{0} from some geodesic loop based at b.

Proof : Writing M = H^{3}/Γ, we note that the convex hull of the Γ-orbit, B, corresponding to b, is the union of all geodesic 3-simplices with vertices in B. Any point in such a 3simplex is a bounded distance from one of its edges. This edge projects to a loop based at b in M. ♢

To prove that e is topologically finite, after passing to the cover corresponding to e, we can suppose that π_{1}(M) is supported on e. We are thus in the situation described in Section 4.

Let ϕ : K ⟶ M be the map described there, and write Φ = ϕ(K). Let Υ be the component of Π(Φ) containing e. Given x ∈ Υ or Q ⊆ Υ we write D(x) = ρ(x, ∂Υ) and D(Q) = ρ(Q, ∂Υ), where ρ is the electric pseudometric described in Section 3.4. Thus D : Υ ⟶ [0, ∞) is a proper function, which we can think of as measuring “depth” in the end.

Let a be some fixed vertex of K. Since e is degenerate, we can find a sequence of points, x_{i} ∈ core(M) with x_{i} → e. In other words, D(x_{i}) → ∞. Moreover, we can find properly embedded rays τ_{i} based at x_{i}, with D(τ_{i}) → ∞. By Lemma 5.2, for each i, there is a loop, α_{i}, based at a, such that the corresponding geodesic loop in M based at ϕ(a) passes within a bounded distance of x_{i} (in the hyperbolic metric). We can extend ϕ : K ⟶ M to a map ϕ : K_{α_{i}} ⟶ M as in Section 3.2. For all sufficiently large i, Proposition 4.1 gives us an embedded surface S_{i} in int Υ, of genus g, which separates e from ∂Υ, and which is π_{1}-injective in Π(Φ) and hence also in Υ. Moreover, ρ(τ_{i}, S_{i}) and the ρ-diameter of S_{i} are uniformly bounded. In particular, D(S_{i}) → ∞. In other words, the S_{i} go out the end.

Suppose first that e is incompressible. In this case, each S_{i} is incompressible in M, and so its inclusion into M is a homotopy equivalence. Thus, the S_{i} are all homotopic, and so, by Lemma 2.5.2, e is topologically finite.

We can thus assume that e is compressible. In this case, we show (cf. [Sou]):

Lemma 5.3 : The S_{i} lie in finitely many homotopy classes in Υ.

Proof : Since e is compressible none of the S_{i} are incompressible in M. Thus, by Dehn’s lemma, there is some essential simple closed curve, β_{i} ⊆ S_{i} which is homotopically trivial in M. Now S_{i} is incompressible in Υ, and so (by Lemma 3.2.1 applied to a locally CAT(−1) space), we can find a balanced map ψ_{i} : S_{i} ⟶ Υ homotopic to the inclusion, S_{i} ↪ Υ, with ψ_{i}(β_{i}) a closed (local) geodesic in the intrinsic path-metric on Υ. Since this cannot be geodesic in M, we must have ψ_{i}(β_{i}) ∩ ∂Υ ≠ ∅. In particular, each ψ_{i}(S_{i}) meets ∂Υ.

Now, since Φ is compact and fat, the injectivity radius of Υ on ∂Υ is positive; that is, any loop in Υ that is non-trivial in Υ and meets ∂Υ has length at least some ζ > 0. It follows that any essential loop in Υ of length less that ζ lies in a Margulis tube in M. Using Lemma 3.4.1, we conclude that the ρ-diameters of the sets ψ_{i}(S_{i}) are bounded. Since the function D is proper, they all lie in some compact subset R ⊆ Υ. Let η > 0 be the injectivity radius of Υ on R. If ψ_{i} and ψ_{j} agree everywhere to within a distance η, they must be homotopic in Υ, by linear homotopy along short geodesic segments. It now follows easily that there are only finitely many possibilities for the homotopy class, as claimed. ♢

Thus, passing to a subsequence, we can assume that the S_{i} are all homotopic in Υ and so e is topologically finite, again by Lemma 2.5.2.

This proves Proposition 5.1.

Now, as noted in Section 2.3, M has only finitely many ends, and it follows that M is topologically finite, proving Theorem 0.1 in the case without cusps.

6. Tameness with cusps.

6.1. Outline.

In this section, we explain how to adapt the earlier arguments to give a proof of topological finiteness in the general case (Theorem 0.1). We therefore have to deal with Z-cusps and Z ⊕ Z-cusps, but only the former cause significant complications. Much of the modification is fairly routine, though there are some more subtle points that call for more detailed comment. We work through each section in turn.

6.2. Topological constructions in the relative case.

Typically we will be dealing with a pair of spaces, (A, B), where A is connected, and B ⊆ A is closed. We often loosely refer to B as the “peripheral structure” since it will be associated to cusps or boundary components. Our notation will frequently suppress explicit reference to B. We shall assume that A an B are reasonably nice spaces (in practice manifolds or simplicial complexes).

The fundamental group π_{1}(A) carries a peripheral structure namely a set of conjugacy classes of subgroups — those supported on the connected components of B.

A standard trick for dealing with the situation when B ≠ ∅ is to take the double DA = D_{B}A, i.e. take two disjoint copies of A and glue them along B. Note that π_{1}(DA) is represented by a graph of groups with two vertex groups isomorphic to π_{1}(A) connected by a set of edges, one for each (conjugacy class of) peripheral subgroup.

Suppose that C ⊆ A is closed and meets every component of B. Then DC ⊆ DA is connected. If π_{1}(C) ⟶ π_{1}(A) is surjective then it is not hard to verify that π_{1}(DC) ⟶ π_{1}(DA) is also surjective.

We now move on to consider more specific cases. Suppose that S is a compact surface, neither a disc nor an annulus, with boundary ∂S, which we view as its peripheral structure. (Here, the letter Σ will be reserved for surfaces without boundary, compact or otherwise.) The type of S is the pair (g, p) where g is the genus, and p the number of boundary components. A convenient measure of the complexity of S will be its first Betti number, b(S) = dim H_{1}(S) = 2g + (p − 1)_{+} where x_{+} = max(0, x). If p > 0 then b(S) is also the genus of the closed surface DS. In fact, b(DS) = 2b(S).

Suppose that F is a disjoint union of non-annular subsurfaces of a closed surface Σ, and that no boundary component of F is homotopically trivial in Σ. One can check that b(F) ≤ b(Σ). (Of course, the map H_{1}(F) ⟶ H_{1}(Σ) need not be injective in general.)

Returning to S, we can form the product manifold, P = S × [0, 1]. We refer to ∂^{V} P = ∂S × [0, 1] and ∂^{H}P = S × {0, 1} as the vertical and horizontal boundaries respectively. (This ties in with the terminology used in [Mi,Bow2,Bow3] etc.) We think of ∂^{V} P as the peripheral structure of P. (Thus, DP ≡ DS × [0, 1].)

We can define a “relative compression body” by the following construction. Take a (possibly empty) collection, P_{i} = S_{i}×[0, 1], of product manifolds, for i = 1, …, n, and take an embedded disc, D_{i} ⊆ (int S_{i}) × {1}. Now take a handlebody, H, of genus g_{0} (possibly 0), and a set, D′_{1}, …, D′_{n}, of disjoint embedded discs in ∂H. We identify D_{i} with D′_{i} to give us a connected manifold, P. Let ∂^{V} P = ⋃_{i=1}^{n} ∂^{V} P_{i}, and let ∂^{H}P = ∂P \ int ∂^{V} P. Now, ∂^{H}P consists of a set of inner boundaries S_{i} × {0}, and an outer boundary component S = ∂_{0}^{H}P. If S_{i} has type (g_{i}, p_{i}), then S has type (g, p), where g = g_{0} + g_{1} + ⋯ + g_{n} and p = p_{1} + ⋯ + p_{n}. We refer to (g_{0}, g_{1}, …, g_{n}, p_{1}, …, p_{n}) as the type of P. A manifold of this sort, with peripheral structure, ∂^{V} P, is referred to as a relative compression body. (This ties in with the inductive definition of compression body described in Section 2.2.)

We shall refer to a group, G, if the from π_{1}(P), with its peripheral structure as a relative compression group, of genus g, and complexity b(G) = b(∂_{0}^{H}P) = 2g + (p − 1)_{+}, where p is the number of conjugacy classes of peripheral subgroups. (Note that b(G) is not the same as dim H_{1}(P).) We see that π_{1}(P) = F_{g_{0}} ∗ (∗_{i}π_{1}(S_{i})), with p_{i} peripheral subgroups conjugate into π_{1}(S_{i}).

Note that DP is a compression body of type (2g_{0}, g_{1}, g_{1}, g_{2}, g_{2}, …, g_{m}, g_{m}, g′_{m+1}, g′_{m+2}, …, g′_{n}), where we assume that the p_{i} = 0 if and only if i ≤ m, and set g′_{i} = 2g_{i} + p_{i} − 1 for i > m. From the canonical structure of π_{1}(DP) as a free product of surface groups and the free group F_{2g_{0}}, and counting the number of peripheral subgroups in each of the surface group factors, we see that this type, and hence g_{0}, (g_{1}, p_{1}), …, (g_{n}, p_{n}) are determined by the structure of π_{1}(P). In other words, the type of a relative compression group is well defined (up to consistently permuting the g_{i} and p_{i}). Moreover, the factors π_{1}(S_{i}) are well defined up to conjugacy in π_{1}(P). We refer to these as the surface factors.

We have the following analogue of Lemmas 2.2.1 and 2.2.2.

Lemma 6.2.1 : Suppose that P is a compact manifold and that F ⊆ ∂P is a disjoint union of closed annuli, pairwise non-homotopic in ∂P. Suppose that π_{1}(P) is a relative compression group, with peripheral structure coming from F. Suppose that each of the surface factors of π_{1}(P) is conjugate into the fundamental group of a component of ∂P \F. Then P is a relative compression body with ∂^{V} P = F.

Lemma 6.2.2 : Suppose that P is a compact manifold and that F ⊆ ∂P is a disjoint union of closed annuli. Suppose that S is a component of ∂P \ int F, and that π_{1}(S) ⟶ π_{1}(P) is surjective. Then P is a relative compression body with vertical boundary ∂^{V} P = F and outer boundary ∂_{0}^{H}P = S.

These follow by similar arguments, decomposing along compression discs. We use the fact that if Q is a compact manifold and S ⊆ ∂Q is a closed subsurface with π_{1}(S) ⟶ π_{1}(Q) an isomorphism, then Q ≅ S × [0, 1] with S ≡ S × {0} (see Theorem 10.2 of [H]).

Let R be a (non-compact) 3-manifold with boundary. As usual we assume our manifolds to be aspherical. The discussion of Section 2.3 goes through much as before, with R in place of M. Typically we use relative second homology, H_{2}(P, ∂^{V} P) or H_{2}(R, ∂R), but the usual (non-relative) first homology, H_{1}(P) or H_{1}(R).

The following strengthening of the Scott core theorem is due to McCullough [Mc].

Theorem 6.2.3 : [Mc] Suppose R is a 3-manifold with π_{1}(R) finitely generated, and that F ⊆ ∂R is a compact subsurface. Then there is some compact submanifold, P ⊆ R, with P ↪ R a homotopy equivalence, and with P ∩ ∂R = F. ♢

Suppose that ∂R is a union of bi-infinite cylinders, and that there are no proper essential discs or annuli. Let F ⊆ ∂R be union of compact cores of a finite set of these cylinders, and let P be as given by Theorem 6.2.3, so that F ⊆ ∂P. Now the Euler characteristic of ∂P is twice the Euler characteristic of P, which is determined by π_{1}(M). Thus, there is a bound on the number of non-toroidal components of ∂P and hence the number of components of F (since no two of them can be homotopic). We can deduce that ∂R, in fact, has only finitely many components, and so we can retrospectively assume that F contains a core from each such component. (This is a standard topological proof of Sullivan’s cusp finiteness theorem [Su], see [FeM].) In this situation, we will write ∂^{V} P = F.

We say that an embedded surface S in a 3-manifold, R, is proper if ∂S = S ∩ ∂R. Suppose now that ∂R is incompressible and anannular (that is, any proper disc or annulus in R can be homotoped into ∂R). Suppose that ∂R is a union of bi-infinite cylinders. Let e be an isolated end of R. It has a one-ended neighbourhood, E, with E ∩ ∂D consisting of semi-infinite cylinders, S^{1} × [0, ∞). We write ∂^{V} E = E ∩ ∂R, and write ∂^{H}E for the relative boundary of E in M. Thus, ∂^{H}R is a proper surface in R. We write p(e) for the number of components of ∂^{V} E. (This is independent of E.) As before, we say that E is full if π_{1}(R \ E) ⟶ π_{1}(R) is surjective.

The earlier discussion now goes through. The end e has associated with it an end group, G(e), which is a relative compression group. This determines the type of the end e. In particular, we see that it has associated to it a genus, g, and a complexity b(e) = b(G(e)). Note that b(e) = 2g + (p(e) − 1)_{+}.

The above can be arrived at by similar arguments, or assisted by the observation that, if p > 0, then e determines an end, De of DR. Moreover, if E is a full neighbourhood of e in R, then DE is a full neighbourhood of De in DR.

The following result could be elaborated upon (taking account of genus) but will suffice for our purposes. It is the analogue of Corollary 2.4.5, and can be proven similarly or by doubling.

Lemma 6.2.4 : Suppose that P ⊆ R is a closed subset carrying π_{1}(R) and separating the end e from any other end. Suppose that S ⊆ M \ P is a proper surface separating P from e. Then b(S) ≥ b(e). Moreover, if b(S) = b(e), then S is π_{1}-injective in M \ P. ♢

The relative version Waldhausen’s cobordism theorem states that if two proper embedded subsurfaces, S_{1}, S_{2} ⊆ R are homotopic relative to ∂R, then they bound a product region S_{i} × [0, 1].

We now move on to Section 2.6. We say that a map, f : S ⟶ R is proper if f^{−1}∂R = ∂S. For the analogue of Theorem 2.6.1, homotopies are interpreted relative to ∂R. It is discussed in [Bon1]. Here is a generalisation of Theorem 2.6.2:

Theorem 6.2.5 : Suppose that f : S ⟶ R is a proper map inducing injections H_{1}(S) ⟶ H_{1}(R), H_{1}(∂S) ⟶ H_{1}(∂R) and H_{2}(S, ∂S) ⟶ H_{2}(R, ∂R). Given any open neighbourhood, U, of f(S), there is a proper embedding S ↪ U of another surface, S′, with b(S′) ≤ b(S), and with H_{2}(S′, ∂S′) ⟶ H_{2}(R, ∂R) injective.

Proof : We follow the proof of Theorem 2.6.2. At the top of the tower we have maps h : S ⟶ N and λ : N ⟶ U. The map H_{1}(S) ⟶ H_{1}(N) is an isomorphism, and so b(∂N) = dim H_{1}(∂N) ≤ 2 dim H_{1}(N) = 2b(S). Let ∂^{V} N = λ^{−1}(∂R) and let ∂^{H}N be the closure of ∂N \ ∂^{V} N. Note that ∂^{V} N is a disjoint union of essential annuli. We decompose ∂^{H}N as A ⊔ B, so that h(S) homologically separates A and B (in terms of relative homology, H_{2}(S, ∂S) and H_{2}(N, ∂^{V} N)). Now the boundary components of ∂^{H}N are all essential in the closed surface ∂N. From an earlier observation, we see that b(A) + b(B) = b(A ∪ B) ≤ b(∂N) ≤ 2b(S). We can thus assume that b(A) ≤ b(S).

We can now continue as before. On going down the tower, we have the additional possibility of having to carry out surgery on a double arc connecting boundary components (in addition to double curves). However, as before, this does not change the Euler characteristic, and so the first Betti number also remains unchanged. At the bottom of the tower, we arrive at our desired surface. ♢

Moving on to Section 2.7, we have the following generalisation of Lemma 2.7.1, which is proven in [Bow3].

Lemma 6.2.6 : Suppose that R is a 3-manifold and S ⊆ ∂R is a compact subsurface and that no component of ∂S is homotopically trivial in S. Suppose that S can be homotoped into a closed subset, P ⊆ R with P ∩ S = ∅. Then S is π_{1}-injective in R \ P. ♢

(The homotopy referred to need not be relative to ∂R, though this case would be sufficient for our application here.)

We have the following slight strengthening of Theorem 2.8.1.

Theorem 6.2.7 : Let N be an (aspherical) topologically finite 3-manifold (without boundary), such that each π_{1}-injective torus in N can be homotoped out an end of N. We suppose that N has at least one non-toroidal end (i.e. genus at least 2). Let X be a cover of N with π_{1}(X) finitely generated. Then X is topologically finite.

Proof : The argument is again due to Thurston. By Thurston’s hyperbolisation theorem [O,K], N is homeomorphic to H^{3}/Γ where Γ = π_{1}(N) is a geometrically finite kleinian group acting on H^{3} with no Z-cusps. Each non-toroidal boundary component of the compactification of N can be identified with a component of Ω(Γ)/Γ where Ω(Γ) is the discontinuity domain. Let Y ⊆ H^{3} be the closed 1-neighbourhood of the convex hull of the limit set, Λ(Γ), of Γ, and let B be a (possibly empty) strictly invariant collection of horoballs (coming from the Z ⊕ Z-cusps). We write B for the interior of ⋃ B. Thus, (Y \ B)/Γ is compact, and Y \ B ⊆ N(∂Y \ B, r) for some r ≥ 0. Let G ≤ Γ correspond to π_{1}(X) ⊆ π_{1}(N), and let Z be the 1-neighbourhood of the convex hull of Λ(G). Thus Z \ B ⊆ N(∂Z \ B, r). By Ahlfors’s Finiteness Theorem [Ah], (∂Z(G) \ B)/G is compact, and so (Z \ B)/G is compact. It follows that G is geometrically finite. (Note that if we only remove those horoballs with non-trivial G-stabilisers, it is easily seen that the above quotient remains compact, and this is one of the standard characterisations of geometrical finiteness for kleinian groups.) In particular, int Z/G ≅ H^{3}/G ≅ X is topologically finite.

♢

There is an addendum to the above result. Suppose that N̄ is a manifold compactification of N and that N̂ is a manifold with N ⊆ N̂ ⊆ N̄. Lifting, we see from the above argument, that we can embed the cover, X̂ of N̂ in the compactification of X̄. In our application of this, each boundary component of N̂ is a cylinder which lifts to a cylinder in X̂. In fact, this addendum can be bypassed, but only at the cost of introducing another, more complicated, argument, as we discuss at the relevant time (after Lemma 6.4.1).

6.3. Polyhedra with cusps.

The main novelty here is that we allow our hyperbolic simplices to have ideal points. We can extend continuously over “blow-ups” around our ideal points. Rather than attempt to give a formal unified account, we deal separately with manifolds and complexes.

Let M be a complete hyperbolic manifold. The non-cuspidal part, Ψ(M), of M is obtained by removing the interiors of all Margulis cusps from M. The thick part, Θ(M) ⊆ Ψ(M), is then obtained by removing Margulis tubes from Ψ(M). In what follows, we shall deem a certain set of Z-cusps of M be be essential, and construct a manifold, R(M) by removing only the essential cusps. Thus Ψ(M) ⊆ R(M) ⊆ M. (All the Z ⊕ Z-cusps are considered inessential.) Note that each component of ∂R(M) is a bi-infinite cylinder. We can write M \ int R(M) as ∂R(M) × [0, ∞), where ∂R(M) ≡ ∂R(M) × {0}, and each {x} × [0, ∞) is a geodesic ray. We can adjoin a copy of ∂R(M) to M as ∂R(M) × {∞}. The resulting manifold, denoted M ∪ ∂^{V} M, is homeomorphic to R. It has boundary ∂^{V} M ≅ ∂R, which we view as the peripheral structure.

We can do something similar with simplicial complexes. Suppose that K̄ is a locally finite simplicial complex, and that A ⊆ V (K̄) is a set of vertices. Let K = K̄ \ A. We define the blow-up of, K ∪∂^{V} K, by adjoining a copy of the link of a, for each vertex a ∈ A.

(Note that K ∪ ∂^{V} K can also be triangulated as a simplicial complex.)

This construction can be given a geometric interpretation. Suppose that K is built out of hyperbolic simplices in such a way that each “missing” vertex in A corresponds to an ideal point. Then the points of ∂^{V} K can be thought of as obtained by adjoining an ideal point to each geodesic ray.

We can now generalise the notion of a polyhedron in M. This is a simplicial complex, Φ ⊆ M, where the simplices are allowed to have ideal vertices in the essential cusps of M. We see that the inclusion Φ ↪ M extends naturally to an injective map Φ ∪ ∂^{V} Φ ↪ M ∪ ∂^{V} M.

More generally, we have an obvious notion of polyhedral map, ϕ : K ⟶ M, where again, ideal vertices go to essential cusps. This has a natural extension, ϕ : K ∪ ∂^{V} K ⟶ M ∪ ∂^{V} M. Moreover, Φ = ϕ(K) is a polyhedral complex, with Φ ∪ ∂^{V} Φ = ϕ(K ∪ ∂^{V} K).

The discussion of fat and balanced polyhedra and maps in Sections 3.1 and 3.2 goes through with little change. When realising maps of finite complexes (cf. Lemma 3.1.1) we can talk about proper homotopy classes of proper maps, K ⟶ M, where the ends of K (corresponding to missing vertices) go out essential cusps of M. This is equivalent to considering relative homotopy classes of maps K ∪ ∂^{V} K, ∂^{V} K ⟶ M ∪ ∂^{V} M, ∂^{V} M.

There is again little change to Section 3.3.

Moving on to Section 3.4, we let d_{R} be the induced path-metric on R(M), and let ρ_{R} be the path-pseudometric obtained from d_{R} by deeming each Margulis tube and each inessential Margulis cusp to have diameter 0.

Now, Σ, becomes a topologically finite surface which can be compactified to a surface Σ ∪ ∂^{V} Σ with boundary ∂^{V} Σ. The metric, σ, on S will be a finite-area singular hyperbolic metric. In Lemma 3.4.1, we get instead, a bound either on the ρ_{R}-diameter of f(Σ)∩R(M), or on ρ_{R}(x, f(β)).

6.4. Wrapping with cusps.

Let M be a complete hyperbolic 3-manifold with π_{1}(M) finitely generated. We suppose, as before, that M has a preferred set of Z-cusps deemed “essential”. Let e be an end of M ∪ ∂^{V} M. We let ∂_{e}^{V} M be the set of components of ∂^{V} M that meet each neighbourhood of e. It is a consequence of Theorem 6.2.3 and the subsequent discussion that we can choose such a neighbourhood so that it intersects each such component in a bi-infinite cylinder. The subgroup, G(e) of π_{1}(M) supported on e is a relative compression group (see Section 6.2). Let N̄(e) be the the associated cover of M ∪ ∂^{V} M, and write N(e) for its interior. The cusps associated to the end e (namely bounded by ∂_{e}^{V} M) lift to N(e). We view these cusps as the essential cusps of N(e). In other words, we can identify ∂^{V} N(e) with ∂_{e}^{V} (M). The manifold N(e) ∪ ∂^{V} N(e) has an “outer” end, e_{0}, which is a lift of e, together with a (possibly empty) set of inner ends.

In general these ends may contain other Z-cusps (traditionally termed “accidental cusps”). Note, however, that any such Z-cusp in e_{0} must have already existed (as an inessential cusp) in M.

To do the construction properly, we therefore begin by deeming all Z-cusps in M to be essential. We can take any end, e, of M ∪ ∂^{V} M, and construct N(e) ∪ ∂^{V} N(e). Then the outer end has no Z-cusps. Replacing M with N(e). We are therefore reduced to the following situation:

(∗) M is a complete hyperbolic 3-manifold with an associated set of “essential” Z-cusps. The manifold M ∪ ∂^{V} M has an “outer” end e, which meets each component of ∂^{V} M in an unbounded set. The fundamental group, π_{1}(M), is supported on (any neighbourhood of) e. The end e contains no Z-cusps.

Now π_{1}(M) is a relative compression group, and M ∪∂^{V} M has a (possibly empty) set of “inner” ends e_{1}, …, e_{n}. Such an inner end may contain Z-cusps, or may be a Z⊕Z-cusp. If e is itself a Z ⊕ Z-cusp, then it is clearly topologically finite, so we can ignore this case.

We will again give separate attention to the cases where e is compressible or incompressible. In the former case, we will assume all the inner ends to be topologically finite. In the latter case, we will not be able to assume that, but much of the reasoning is simpler anyway. The following discussion deals mostly with the compressible case.

Let K_{0} be a compact 2-complex with a balanced polyhedral homotopy equivalence, Φ : K_{0} ⟶ M. Each essential cusp of M corresponds to a closed curve in K which we take to lie in the 1-skeleton. We can form another compact complex by coning over each such curve. Removing the cone points we then obtain a complex, K, with “missing vertices”. We can now extend ϕ to a proper polyhedral homotopy equivalence ϕ : K ⟶ M, where the ends of K get sent to the essential cusps. This extends further to a continuous map ϕ : K ∪ ∂^{V} K ⟶ M ∪ ∂^{V} M, which is again a homotopy equivalence. (Note that ∂^{V} K is a union circles.) Let Φ = ϕ(K). We similarly define K_{α}, and let Φ_{α} = ϕ(K_{α}). Now Φ_{α} ∪ ∂^{V} Φ_{α} separates the ends of M ∪ ∂^{V} M, and we let Υ ∪ ∂^{V} Υ be the closure of the complementary component containing e. (Here Υ ⊆ M and ∂^{V} Υ ⊆ ∂^{V} M.) We write ∂^{H}Υ = (Υ ∪ ∂^{V} Υ) ∩ (Φ ∪ ∂^{V} Φ). This is the relative boundary of Υ ∪ ∂^{V} Υ in M ∪ ∂^{V} M.

In this case, we will assume that our ray τ lies in Υ ∩ R(M), based in Φ_{α} ∩ R(M), and going out the end. We will assume that ρ_{R}(∂^{H}Υ ∩ R(M), τ) is sufficiently large.

We now define W ∪ ∂^{V} W ⊆ M ∪ ∂^{V} M with relative boundary ∂^{H}W, similarly as we did with W ⊆ M in the case where ∂^{V} M = ∅. This manifold contains all the inner ends of M ∪ ∂^{V} M. We have the following version of the statement that Δ is atoroidal (cf. Lemma 4.2).

Lemma 6.4.1 : Any π_{1}-injective torus in W can be homotoped out an end of N.

Proof : Let Δ be an embedded π_{1}-injective torus. If Δ is π_{1}-injective in M, then it bounds a Z ⊕ Z-cusp. This cannot contain ∂^{H}W and so Δ bounds some cusp in W. We can therefore assume that Δ is not π_{1}-injective in M, and the argument proceeds as with Lemma 4.2. ♢

As before, we can now take a cover, X, of W so that ϕ : K ⟶ W lifts to a homotopy equivalence, ϕ̃ ⟶ X. By Lemma 6.2.7, X is topologically finite. In other words we can embed it in a compact manifold X̄. By the addendum to Lemma 6.2.7, we can lift ∂^{V} W to a subset ∂^{V} X ⊆ ∂X̄, so that the closure of ∂^{V} X is a disjoint union of annuli. (The addendum can be bypassed using the observation that a properly embedded essential semiinfinite cylinder, S^{1}×[0, ∞) cannot be knotted in a 3-manifold. However, this requires a bit of work to prove.) By Lemma 6.2.1, X̄ is a relative compression body with respect to this peripheral structure. We write ∂^{H}X ⊆ X for the outer boundary. Thus, X ∪ ∂^{V} X ∪ ∂^{H}X is a relative compression body with the inner boundary components removed. Pushing the surface ∂^{H}X slightly inside this manifold, while keeping its boundary components in ∂^{V} X, we a surface, denoted ∂^{H}P, in X ∪ ∂^{V} X. This is the outer boundary of another relative compression body, P ∪ ∂^{V} P ∪ ∂^{H}P, with P ⊆ X and ∂^{V} P ⊆ ∂^{V} X.

We now proceed to construct a locally CAT(−1) space Z pretty much as before. It has a vertical boundary, ∂^{V} Z, coming from ∂^{V} X. There is a natural map λ : Z ⟶ M, which extends to a relative homotopy equivalence Z ∪ ∂^{V} Z ⟶ M ∪ ∂^{V} M.

Next, we construct maps h, f : Σ ⟶ Υ as before, by homotoping ∂^{H}P in Z ∪ ∂^{V} Z. By construction, the maps are just products in some neighbourhood of the essential cusps (i.e. they send rays to geodesic rays). In fact, we can take these cusps to be uniform. By construction, we can then assume that the 1-skeleton of the triangulation does not enter the cusps. Also by construction, Σ has the same type as the outer boundary component of the relative compression body. In particular (from the definition of the complexity of an end) we have b(Σ) = b(e). Now f(Σ ∪ ∂^{V} Σ) is a properly embedded surface that bounds a manifold Q ∪ ∂^{V} Q ⊆ Z ∪ ∂^{V} Z. We write ∂^{H}Q = f(Σ ∪ ∂^{V} Σ) for the relative boundary. Again Q ∪ ∂^{V} Q ∪ ∂^{H}Q is a relative compression body with the inner boundary components removed. The map λ : Q ∪ ∂^{V} Q, ∂^{V} Q ⟶ M ∪ ∂^{V} M, ∂^{V} M is a relative homotopy equivalence.

The analogue of Lemma 4.7 now follows by applying the analogous statement to Lemma 4.5, where we need to take care that the common image of the curves α and β is not homotopic into Z ⊕ Z-cusp of M.

We now have:

Lemma 6.4.2 : f : Σ ⟶ Υ induces an injection H_{2}(Σ ∪ ∂^{V} Σ, ∂^{V} Σ) ⟶ H_{2}(Υ ∪ ∂^{V} Υ, ∂^{V} Υ).

Proof : The case where M is a handlebody (with ∂^{V} M = ∅) has already been accounted for (Lemma 4.7).

In all other cases, f induces an injection to H_{2}(M ∪ ∂^{V} M, ∂^{V} M) as with Lemma 4.7, and the result follows. ♢

Note that, by construction, the map H_{1}(∂^{V} Σ) ⟶ H_{1}(∂^{V} Υ) is injective.

Lemma 6.4.3 : f : Σ ⟶ Υ induces an injection H_{1}(Σ) ⟶ H_{1}(Υ).

Proof : This is most conveniently seen via the doubling trick. We get doubled maps Dϕ̃ : DK ⟶ DQ and Dλ : DQ ⟶ DM, which are both homotopy equivalences. We argument proceeds as before to show that H_{1}(DΣ) ⟶ H_{1}(DΥ) is injective. But H_{1}(Σ) ⟶ H_{1}(DΣ) is injective, and the result follows. ♢

Using Theorem 6.2.5, we can find a proper embedded surface F ⊆ Υ ∪ ∂^{V} Υ, which is non-trivial H_{2}(Υ ∪ ∂^{V} Υ, ∂^{V} Υ), and with b(S) ≤ b(e). In particular, we can assume that S ∩ R(M) is in a small neighbourhood f(Σ) ∩ R(M). Thus, S ∩ R(M) has bounded ρ_{R}- diameter, and is a bounded ρ-distance from τ ⊆ Υ. (Since we have a bound on complexity, the bounded diameter lemma applies.) After pushing contractible components of S \R(M) into M, we we can assume that each component of S\R(M) is an annulus (and so S∩R(M) is homeomorphic to S). Indeed, we can modify S so that it is a product in a neighbourhood of the cusps (i.e. is ruled by geodesic rays going out the cusps). Thus, S ⊆ Υ. Applying Lemma 6.2.4, we see that b(S) = b(e).

In summary, we have found an embedded π_{1}-injective surface, S ⊆ Υ, of bounded complexity, with S ∩ R(M) of bounded ρ_{R}-diameter, and with ρ(τ, S) bounded.

The argument in the case where e is compressible requires the same modification as in the case with no cusps.

6.5. Proof of tameness when there are cusps.

We begin with a fairly general and straightforward observation. Suppose M is a complete manifold with π_{1}(M) finitely generated, and with an associated set of essential Z-cusps. Let e be an end M ∪ ∂^{V} M. This end may be subdivided by inessential Z-cusps in e. (That is, several end of M minus the inessential cusps might lie in a single end of M.) But if each of the subdivided ends is topologically finite, then so is e.

To prove that an end of M∪∂^{V} M is topologically finite, we proceed by induction on the complexity, b(e), of e. By the discussion at the beginning of Section 6.4, we can reduce to the case satisfying (∗). In this case, one easily checks that either e is incompressible, or else all of the inner ends have strictly lower complexity and hence can be assumed topologically finite. (They may be subdivided by inessential Z-cusps, but that can only lower complexity further.) We start the induction with the case of tori (which are necessarily Z ⊕ Z-cusps) or 3-holed spheres (which are necessarily geometrically finite).

We now obtain a sequence, S_{i}, of properly embedded surfaces in M ∪ ∂^{V} M going out the end, of bounded complexity, b(S).

We need to check:

Lemma 6.5.1 : The S_{i} lie in finitely many homotopy classes in Υ ∩ R(M) (relative to Υ ∩ ∂R(M)).

Proof : The argument proceeds much as with Lemma 5.3. We assume that e is compressible. Since b(S_{i}) is bounded, we can assume that each S_{i} is the image of an embedding of a fixed surface, S. We can find balanced maps ψ_{i} : S ⟶ Υ ∪ ∂^{V} Υ in the same homotopy classes as the respective embeddings, so that ψ_{i}(β_{i}) is geodesic in Υ, where β_{i} is a simple closed curve in S with ψ_{i}(β_{i}) homotopically trivial in M. Thus, ψ_{i}(β_{i}) ∩ ∂Υ ≠ ∅. If the intersections ψ_{i}(β_{i}) ∩ ∂Υ all lie in a compact set, we are done, by essentially the same argument. However, a-priori ψ_{i}(β_{i}) might meet ∂Υ only far out the cusps. This is prevented by the observation that, in the induced metric on int S_{i}, β_{i} is also geodesic and (homotopic to) a simple closed curve. Since S_{i} is singular hyperbolic, it is easily verified that there is a bound on how far β_{i} can go out a cusp. ♢

Thus, after passing to a subsequence, we can assume that the S_{i} are all homotopic in Υ ∩ R(M), and it follows by the relative version of Waldhausen’s cobordism theorem, that e is topologically finite.

Finally, using the relative Scott theorem [Mc] it follows that R(M) has only finitely many ends and so R(M) hence M is toplogically finite. In fact, the argument has shown directly that the non-cuspidal part, Ψ(M), is M is topologically finite, in the sense that we can embed Ψ(M) in a compact manifold M′, so that the interiors of Φ(M) and M′ coincide, and such that the boundary of Ψ(M) is a subsurface of ∂M′.

6.6. Orbifolds.

We finally note that the tameness theorem extends to orbifolds in the following sense. Any complete hyperbolic orbifold, O, with orbifold fundamental group Γ is the quotient of hyperbolic 3-space by a properly discontinuous action of Γ. If Γ is finitely generated, then Selberg’s Lemma tells us that it is virtually torsion free, and so O is the quotient of a hyperbolic manifold, M, by the action of a finite group, G. Tameness tells us that M is homeomorphic to the interior of a compact manifold M′, and Theorem 8.5 of [MeS] tells us that we can take this homeomorphism to be equivariant with respect to an action of G on M′. It follows that O is topologically finite, in the sense that it is the interior of a compact orbifold (with boundary), M′/G.

7. Variable curvature.

In this section, we describe how the main ideas go through, with some modification, to the case of pinched negative curvature. Our goal is to prove the following (earlier stated as Theorem 0.4).

Theorem 7.1 : Let M be a complete riemannian 3-manifold of pinched negative curvature with π_{1}(M) finitely generated. Then M is diffeomorphic to the interior of a compact manifold with boundary.

By “pinched negative curvature” we mean that all sectional curvatures lie between two negative constants. Note that we can write M = Ξ/Γ where Ξ is a pinched Hadamard manifold (i.e. simply connected), and Γ ≅ π_{1}(M). Thus Ξ plays the role of H^{3} in what follows.

We remark that we also get tameness of orbifolds, as discussed in Section 6.6. However, in this more general situation, we do not have Selberg’s Lemma, so we need to make the assumption that the orbifold fundamental group is virtually torsion free. In retrospect, one can show that a topologically finite negatively curved orbifold is finitely covered by a manifold, using Thurston’s orbifold theorem (see [BoiLP] or [CoHK]). Thus, one can view tameness as equivalent to being virtually torsion free. I don’t know whether is automatically implied by finite generation. In what follows we will deal only with (orientable) manifolds.

Many aspects of the discussion go through with little change from constant curvature. For example, M has a thick-thin decomposition, where the thin part consists of Margulis tubes and cusps, which (at least in dimension 3) are smooth submanifolds. (The boundaries need not be intrinsically euclidean, but that does not matter here.)

The issue of convex hulls and convex cores is a bit more subtle, since they can no longer be constructed out of simplices. However, the construction of Anderson [An] as elaborated on in [Bow1] gives us what we need here. In particular, we have the following.

If Q ⊆ Ξ is closed, we write join(Q) for the union of all geodesic segments connecting pairs of points of Q, and hull(Q) for the closed convex hull of Q (i.e. the closure of the convex hull).

Lemma 7.2 : There is some r > 0, depending only on the pinching constants, such that for any closed Q ⊆ Ξ, we have hull(Q) ⊆ N(join(Q), r).

From this, we get immediately the generalisation of Lemma 5.2:

Lemma 7.3 : Suppose M is a complete riemannian 3-manifold of pinched negative curvature, and b ∈ M. Then any point of core(M) lies a distance at most r from some geodesic loop based at b, where r > 0 depends only on the pinching constants. ♢

(In fact, this holds in any dimension, if we allow r to depend on dimension as well.) Here core(M) denotes the closed convex core.

We can give the same definition of geometrically finite end of M ∪ ∂^{V} M, namely one which has a neighbourhood that does not meed core(M). Such ends are also easily seen to be topologically finite (see [Bow1]). Moreover (again in this dimension) cusps are also easily seen to be topologically finite. We are therefore reduced to understanding the case of non-geometrically finite ends.

The main complication arises in adapting the notion of a polyhedral complex, or polyhedral map. For a 1-complex, the same definition makes sense. However, to extend over 2-simplices, we would need to use ruled surfaces, or something similar, instead of totally geodesic ones. It seems intuitively clear that the theory should go through much as before, with these notions. However, there are some complicated technical issues, so we suggest an alternative means which avoids the worst of these. (In this discussion, we only need a negative upper curvature bound.)

We say that a locally finite embedded graph Φ ⊆ M is polygonal if each edge is a geodesic segment. We have the same notion of balanced as before: the tangent vectors to Φ at any vertex do not lie in any open hemisphere. Let Π̂(Φ) be the completion of the universal cover, of M \ Φ. We write int Π̂(Φ) ⊆ Π̂(Φ) for the universal cover of M \ Φ.

Lemma 7.4 : If Φ ⊆ M is a balanced polygon, then Π̂(M) is (globally) CAT(−1)

Proof : (Sketch) Suppose x ∈ Φ. We can find some small t > 0 so that Φ ∩ N(x, t) just consists of geodesic segments emerging from x. (We can identify ∂N(x, t) with the unit tangent space, Δ_{x}(M) at x via the exponential map.) We claim that ∂N(x, t)\Φ ↪ M \Φ is π_{1}-injective. For otherwise, by Dehn’s lemma, there would be an embedded disc D ⊆ M \ (Φ ∪ int N(x, t)) with ∂D = D ∩ N(x, t) an essential curve in ∂N(x, t) \ Φ. By gluing in a small disc, we can extend this to a sphere in M meeting Φ only in the point x. This sphere bounds a ball, B ⊆ M. By considering the point of Φ ∩ B furthest from x, after lifting to H^{3}, we contradict the fact that Φ is balanced.

This means that the universal cover of N(x, t) \ Φ is embedded in int Π̂(Φ). Since this holds true for any x ∈ Φ, it is not hard to see that Π̂(Φ) is simply connected. This reduces us to showing that Π̂(Φ) is locally CAT(−1). Again from the above discussion this is now a fairly simple exercise using ruled surfaces. ♢

We have the complication that Π̂(Φ) is not usually locally compact. However, since it is a complete, we still have a classification of isometries of Π̂(Φ) into elliptic, parabolic and loxodromic. If Φ is finite, then in π_{1}(M \ Φ) the only elliptics are those that fix an edge of Φ, and the only parabolics are those that already existed in M.

Given a subset Q ⊆ M \ Φ, we write Π̂_{Q}(Φ) for the quotient of Π̂(Φ) corresponding to the image of π_{1}(Q) in π_{1}(M). This is locally CAT(−1). We get a natural 1-lipschitz map from Π̂_{Q}(M) to M. The preimage Φ is a locally finite graph, and we can construct a collar neighbourhood of it in Π̂_{Q}(M), homeomorphic to a surface times an open interval. For this reason, the topological 3-manifold arguments used in Sections 4 and 5 go through much as before.

Suppose now that K is a locally finite simplicial 2-complex, with 1-skeleton K^{1}. We can define a polyhedral map as a proper map ϕ : K ⟶ M such that each edge gets mapped to a geodesic. We generally regard the maps on 2-simplices to be defined only up to homotopy relative to their boundaries. (We could always take these to be ruled surfaces, but with our modified constructions, their geometry becomes irrelevant.) We say that ϕ is balanced if ϕ|K^{1} is balanced in the obvious sense, reintepreting the definition in constant curvature. Note that ϕ(K^{1}) is a balanced polygonal set in M.

The wrapping construction only requires slight modification. We have balanced maps ϕ : K ⟶ M and ϕ : K_{α} ⟶ M. This time, we write Φ = ϕ(K^{1}) and Φ_{α} = ϕ(K_{α}^{1}). We let F ⊆ M \ Φ_{α} be a surface of minimal genus (or complexity) separating e from Φ_{α}. It follows (using the fact that H_{1}(Φ_{α}) ⟶ H_{1}(M) is surjective) that F is incompressible in M \ Φ_{α}. We define W ⊇ Φ_{α} with ∂W = F as before. Since ∂W is incompressible, we can arrange that ϕ(K_{α}) ⊆ W.

We define Υ_{α} = Π̂_{F} (Φ_{α}). This is locally CAT(−1), and F ↪ Υ_{α} is a homotopy equivalence. We now construct Z as before (using our modified definition of Υ_{α}). Thus, Z ⟶ M \ Φ is a covering space. We end up with maps K ↪ K_{α} ⟶ P ↪ X ↪ Z ⟶ M which are all homotopy equivalences, and the argument now proceeds with little modification.

8. Ahlfors’s finiteness theorem.

In this section, we consider the relationship between tameness and Ahlfors’s Finiteness Theorem, and describe a version of the latter applicable to pinched negative curvature (Theorem 8.8). Such connections have also been observed by Agol. Ahlfors’s theorem [Ah] is probably the central finiteness result in the classical theory of finitely generated kleinian groups.

A kleinian group is a group, Γ, acting properly discontinuously on H^{3}. Thus H^{3}/Γ is a hyperbolic orbifold. It is a manifold if and only if Γ is torsion-free. As observed in Section 6, Selberg’s Lemma tells us that any finitely generated kleinian group is virtually torsion-free. It follows that we can reduce to the torsion-free case in constant curvature.

While tameness gives a new perspective on Ahlfors’s theorem in constant curvature, it does not give an independent argument, since the result has already been used in Theorem 2.8.1. The classical proof uses deformation theory, which also features strongly in the proof of hyperbolisation [O,Ka], so there would seem to be little point in trying to circumvent this. The main interest therefore lies in its adaptability to (pinched) variable curvature, which we discuss at the end of this paper. We explain how tameness gives us back Ahlfors’s Finiteness Theorem (Lemma 8.3 and 8.4 together). Reinterpreting for variable curvature, we will get a new result, namely Theorem 8.8. First, we confine our discussion to constant curvature.

Let M = H^{3}/Γ be a hyperbolic 3-manifold, and let Ω = Ω(Γ) ⊆ ∂H^{3} be the discontinuity domain, and write F_{∞} = F_{∞}(Γ) = Ω/Γ. This is a complex surface. We say that such a surface is analytically finite if it is a finite union of finite type Riemann surfaces, i.e. compact surfaces with finite subsets removed.

Ahlfors’s Finiteness Theorem [Ah] states:

Theorem 8.1 : If Γ is a finitely generated group, then F_{∞}(Γ) is analytically finite.

It turns out that each puncture of F_{∞} is associated to a Z-cusp of M. Related to this, is Sullivan’s finiteness theorem [Su]:

Theorem 8.2 : M has finitely many cusps.

In fact, this can be given a purely topological proof. The fact that M has finitely many Z ⊕ Z cusps is a direct consequence of Scott’s Theorem (Theorem 2.3.1)) — M has finitely many ends. The fact that there are finitely many Z-cusp is a consequence of McCullough’s theorem (Theorem 6.2.3) as observed in Section 6.1 (see also [FeM]).

In fact, part of Theorem 8.1 can also be viewed in this way. Note that by Theorem 6.1.3 [Mc], any compact subsurface, F ⊆ F_{∞} can be assumed to lie in the boundary of a Scott core, P, of M ∪ F_{∞}. The complexity of ∂P is bounded in terms of the Euler characteristic of P, and hence in terms of π_{1}(M) ≅ Γ. This tells us that the union of those components of F_{∞} that are not topological discs or annuli is topologically finite. It does not rule out the possibility of there being infinitely many discs or annuli. (Though in the latter case, they would certainly have to lie in finitely many homotopy classes in M.) Moreover, this particular argument says nothing about the analytic type of F_{∞}.

One can give a more geometric interpretation of Ahlfors’s Finiteness Theorem. Let F_{t} be the boundary of the t-neighbourhood of the convex core of M, for t ≥ 0. For t > 0, this is a properly embedded C^{1}-surface in M.

Lemma 8.3 : Let M = H^{3}/Γ be a complete hyperbolic 3-manifold. The following are equivalent:

(1) F_{∞} is analytically finite,

(2) for some t > 0, F_{t} is analytically finite,

(3) for all t > 0, F_{t} is analytically finite, (4) for some t > 0, F_{t} has finite area,

(5) for all t > 0, F_{t} has finite area,

(6) M ∪ ∂^{V} M has finitely many geometrically finite ends, and each geometrically infinite end has a neighbourhood whose intersection with M is contained in the convex core of M.

In (6), ∂^{V} M is the union of the ideal boundaries of the Z-cusps as defined in Section 6. (All Z-cusps are deemed essential here.) Recall that M ∪ ∂^{V} M is homeomorphic to R(M), that is M with its Z-cusps removed.

Proof : The equivalence of (1)–(5) is well known, and arises from the fact that for all t > 0, the “nearest point” projection F_{∞} ⟶ F_{t} is a quasiconformal homeomorphism, and if 0 < t < u < ∞, the projection F_{u} ⟶ F_{t} is bilipschitz. It is easy to see that the ends of any F_{t} are horocyclic constant curvature cusps — and so do not pose any difficulties.

To relate this to (6), we note that the closure of F_{t} in M ∪ ∂^{V} M can be naturally identified with the compactification F_{t} ∪ ∂^{V} F_{t} of F_{t}. This cuts off a closed neighbourhood, E_{t}, of a geometrically finite end in M ∪∂^{V} M. We can naturally compactify E_{t} to a product manifold, (F_{t} ∪ ∂^{V} F_{t}) × [0, ∞], where we identify int F_{t} × {∞} with F_{∞}. Moreover, any geometrically finite end has this form. Condition (6) tells us that a finite set of such ends account for all of F_{∞}, and the equivalence of (6) with (5) follows easily. ♢

In Lemma 8.3 we did not need that π_{1}(M) be finitely generated. In relation to Ahlfors’s Finiteness Theorem, we need:

Lemma 8.4 : Suppose that M = H^{3}/Γ is a complete hyperbolic 3-manifold with π_{1}(M) finitely generated. Let e be a degenerate end of M. Then there is a neighbourhood, E, of e in M ∪ ∂^{V} M with E ∩ M ⊆ core(M).

This depends on the following observations:

Lemma 8.5 : Let Ψ ⊆ H^{3} be a locally finite balanced polyhedron and let L ⊆ ∂H^{3} be the set of accumulation points of Ψ in ∂H^{3}. Then Ψ ⊆ hull(L).

Proof : If not, there is a closed half-space H ⊆ H^{3} with H ∩ hull(L) = ∅ and with Ψ ∩ int H ≠ ∅. It follows that H ∩ Ψ is compact. If x ∈ Ψ ∩ int H, with d(x, ∂H) maximal, then we see easily that Ψ cannot be balanced at x. ♢

Corollary 8.6 : Suppose that Φ ⊆ M is a finite balanced polyhedron in M (allowing ideal vertices in ∂^{V} M). Then Φ ⊆ core(M).

Proof : Lift to H^{3} and apply Lemma 8.5. ♢ Lemma 8.7 : Suppose that Q ∪ ∂^{V} Q ⊆ M ∪ ∂^{V} M. If ∂Q ⊆ ∂^{V} M ∪ core(M), then Q ⊆ core(M).

(Note that ∂Q \ ∂^{V} Q is the relative boundary of Q in M.)

Proof : The closure of any component of M \ core(M) in M ∪ ∂^{V} M is non-compact and it is easily seen that Q cannot contain such a component. ♢

Proof of Lemma 8.4 : By Section 6.4, we have a sequence of separating surfaces, S_{i} ⊆ M ∪ ∂^{V} M going out the end, e. We can assume them to be products inside a (fixed) set of Z-cusps (i.e. geometric cones of the cusp points). Now S_{i} lies in small neighbourhood of f_{i}(Σ), which in turn lies close to h_{i}(Σ), where f_{i}, h_{i} : Σ ⟶ Υ ⊆ M are proper maps and h_{i} is balanced in Υ. Now h_{i}(Σ) goes out the end, e (and is a product in a neighbourhood of the cusp), and so h_{i}(Σ) ∩ ∂Υ = ∅, for all sufficiently large i. Then h_{i}(Σ) is balanced in M. By Corollary 8.6, we see that h_{i}(Σ) ⊆ core(M). We can therefore assume that f_{i}(Σ) and hence S_{i} ∩ M also lie in core(M).

Now each point of Υ must be separated from e by the surface S_{i}. Thus E = ⋃_{i} E_{i} is neighbourhood of e in M ∪ ∂^{V} M, where E_{i} is the region between S_{0} and S_{i}. Now ∂E_{i} ⊆ ∂^{V} M ∪ core(M) and so by Lemma 8.7, E_{i} ⊆ core(M). Thus E ⊆ core(M) as required. ♢

Now it is easy to see that Lemmas 8.3 and 8.4 together imply Ahlfors’s Finiteness Theorem (Theorem 8.1). Of course we have used this theorem (applied to a different group) in order to prove this. The point of the exercise is that the argument is adaptable to variable curvature.

Let M be a complete manifold of pinched negative curvature, and write F_{t} = N(core(M), t) as before. This is C^{1}-surface, and the projection F_{u} ⟶ F_{t} for u > t is C^{1}-lipschitz and distance decreasing. In particular, area(F_{t}) (if finite) is a non-decreasing function of t.

We can now prove the variable curvature version of the Ahlfors Finiteness Theorem (in the torsion-free case) stated as Theorem 0.5 in the introduction:

Theorem 8.8 : Let M be a complete riemannian 3-manifold of pinched negative curvature, and with π_{1}(M) finitely generated. Let t > 0, and let F_{t} be the boundary of the t-neighbourhood of the convex core of M. Then F_{t} has finite area.

It follows (or can be seen from the argument) that the quotient, F_{∞}, of the discontinuity domain is a topologically finite surface, though it is not clear how to formulate analytical finiteness directly in these terms.

Proof of Theorem 8.8 : Most of the proof of Theorem 8.1 generalises unchanged. (In discussion the balanced maps f_{i}, h_{i} : Σ ⟶ Υ, we can extend over 2-simplices using ruled surfaces. In particular, any degenerate end has a neighbourhood in core(M). Moreover, McCullough’s theorem tells us that M ∪ ∂^{V} M has finitely many ends. What remains is to explain why this implies that area(F_{t}) is finite.

Now in [Bow1] it was shown that if M is geometrically finite then the volume of N(core(M), t) is finite for all t. In particular, for any u > t, vol(C(t, u)) is finite, where C(t, u) = N(core(M), u)\int N(core(M), t). In our situation, M need not be geometrically finite. Nevertheless, C(t, u) lies in the union of the geometrically finite ends. Thus the same argument as given in [Bow1] shows in fact that vol(C(t, u)) is finite. But vol(C(t, u)) = ∫_{t}^{u} area F_{v} dv ≥ (u − t) area F_{t}, and so area F_{t} < ∞ as required. ♢

References.

[Ag] I.Agol, Tameness and hyperbolic 3-manifolds : preprint, Chicago (2004).

[Ah] L.V.Ahlfors, Finitely generated Kleinian groups : Amer. J. Math. 86 (1964) 413–429, 87 (1965) 759.

[An] M.Anderson, The Dirichlet problem at infinity for manifolds of negative curvature : J. Differential Geom. 18 (1983) 701–721.

[BoiLP] M.Boileau, B.Leeb, J.Porti, Geometrization of 3-dimensional orbifolds : Ann. of Math. 162 (2005) 195–290.

[Bon1] F.Bonahon, Structures géométriques sur les variétés de dimension 3 et applications : Thèse, Orsay (1985).

[Bon2] F.Bonahon, Bouts des variétés hyperboliques de dimension 3 : Ann. of Math. 124 (1986) 71–158.

[Bow1] B.H.Bowditch, Geometrical finiteness with variable negative curvature : Duke Math. J. 77 (1995) 229–274.

[Bow2] B.H.Bowditch, Geometric models for hyperbolic 3-manifolds : preprint, Southampton (2005).

[Bow3] B.H.Bowditch, End invariants of hyperbolic 3-manfolds : preprint, Southampton (2005).

[BridH] M.R.Bridson, A.Haefliger, Metric spaces of non-positive curvature : Grundlehren der Math. Wiss. No. 319, Springer (1999).

[BrinJS] M.Brin, K.Johannson, P.Scott, Totally peripheral 3-manifolds : Pacific J. Math. 118 (1985) 37–51.

[BrinT] M.G.Brin, T.L.Thickstun, 3-manifolds which are end 1-movable : Memoirs Amer. Math. Soc. No. 81 (1989).

[BrocB] J.F.Brock, K.W.Bromberg, On the density of geometrically finite Kleinian groups : Acta Math. 192 (2004) 33–93.

[BrocCM] J.F.Brock, R.D.Canary, Y.N.Minsky, Classification of Kleinian surface groups II: The ending lamination conjecture : preprint, Brown/Michigan/Yale (2004).

[Brom] K.Bromberg, Projective structures with degenerate holonomy and the Bers density conjecture : Ann. of Math. 166 (2007) 77-93.

[CalG] D.Calegari, D.Gabai, Shrinkwrapping and the taming of hyperbolic 3-manifolds : J. Amer. Math. Soc. 19 (2006) 385–446.

[Can1] R.D.Canary, Ends of hyperbolic 3-manifolds : J. Amer. Math. Soc. 6 (1993) 1–35.

[Can2] R.D.Canary, A covering theorem for hyperbolic 3-manifolds and its applications : Topology 35 (1996) 751–778.

[CanM] R.D.Canary, Y.N.Minsky, On limits of tame hyperbolic 3-manifolds : J. Differential Geom. 43 (1996) 1–41.

[Ch] S.Choi, The PL-methods for hyperbolic 3-manifolds to prove tameness : preprint, Daejeon (2006).

[CoHK] D.Cooper, C.D.Hodgson, S.Kerckhoff, Three-dimensional orbifolds and cone manifolds : MSJ Mem. No. 5 (2000).

[FeM] M.Feighn, D.McCullough, Finiteness conditions for 3-manifolds with boundary : Amer. J. Math. 109 (1987) 1155–1169, and 112 (1990) 41–45.

[FrHS] M.Freedman, J.Hass, P.Scott, Least area incompressible surfaces in 3-manifolds : Invent. Math. 71 (1983) 609–642.

[G] D.Gabai, Foliations and the topology of 3-manifolds : J. Differential Geom. 18 (1983) 445-503.

[H] J.Hempel, 3-manifolds : Ann. of Math. Studies, No. 86, Princeton University Press (1976).

[JR] W.Jaco, J.H.Rubinstein, PL minimal surfaces in 3-manifolds : J. Differential Geom. 27 (1998) 493–524.

[Ka] M.Kapovich, Hyperbolic manifolds and discrete groups : Progress in Mathematics No. 183, Birkhäuser (2001).

[KlS] G.Kleineidam, J.Souto, Algebraic convergence of function groups : Comment Math. Helv. 77 (2002) 244–269.

[Ma] A.Marden, The geometry of finitely generated kleinian groups : Ann. of Math. 99 (1974) 383–462.

[Mc] D.McCullough, Compact submanifolds of 3-manifolds with boundary : Quart. J. Math. 37 (1986) 299-307.

[MeS] W.H.Meeks III, P.Scott, Finite group actions on 3-manifolds : Invent. Math. 86 (1986) 287–346.

[Mi] Y.N.Minsky, The classification of Kleinian surface groups I : Models and bounds : preprint, Stony Brook (2002).

[My] R.Myers, End reductions, fundamental groups, and covering spaces of irreducible open 3-manifolds : Geom. Topol. 9 (2005) 971–990.

Notes on tameness

[NS] H.Namazi, J.Souto, Non-realizability and ending laminations : in preparation.

[O] J.-P.Otal, Thurston’s hyperbolization of Haken manifolds : Surveys in differential geometry, Vol. III, 77–194, International Press (1998).

[P] L.Person, A piecewise linear proof that the singular norm is the Thurston norm : Topology Appl. 51 (1993) 269–289.

[Sc1] P.Scott, Finitely generated 3-manifold groups are finitely presented : J. London Math. Soc. 6 (1973) 437-440.

[Sc2] P.Scott, Compact submanifolds of 3-manifolds : J. London Math. Soc. 7 (1973) 246-250.

[Som] T.Soma, Existence of ruled wrapping in hyperbolic 3-manifolds : Geom. Topol. 10 (2006) 1173–1184.

[Sou] J.Souto, A note on the tameness of hyperbolic 3-manifolds : Topology 44 (2005) 459–474.

[Su] D.Sullivan, A finiteness theorem for cusps : Acta Math. 147 (1981) 289–299.

[Th1] W.P.Thurston, The geometry and topology of 3-manifolds : notes, Princeton (1979).

[Th2] W.P.Thurston, Three-dimensional manifolds, Kleinian groups and hyperbolic geometry : Bull. Amer. Math. Soc. 9 (1982) 357–381.

[Tu] T.W.Tucker, Non-compact 3-manifolds and the missing-boundary problem : Topology 13 (1974) 267–274.

[Wa] F.Waldhausen, On irreducible 3-manifolds which are sufficiently large : Ann. of Math. 87 (1968) 56–88.

[Wh] J.H.C.Whitehead, A certain open manifold whose group is unity : Quart. J. Math. 6 (1935) 268-279.